### v1

Abstract
While large language models (LLMs) excel at handling long-context sequences, they require substantial key-value (KV) caches to store contextual information, which can heavily burden computational efficiency and memory usage. Previous efforts to compress these KV caches primarily focused on reducing memory demands but were limited in enhancing latency. To address this issue, we introduce FastKV, a KV cache compression method designed to enhance latency for long-context sequences. To enhance processing speeds while maintaining accuracy, FastKV adopts a novel Token-Selective Propagation (TSP) approach that retains the full context information in the initial layers of LLMs and selectively propagates only a portion of this information in deeper layers even in the prefill stage. Additionally, FastKV incorporates grouped-query attention (GQA)-aware KV cache compression to exploit the advantages of GQA in both memory and computational efficiency. Our experimental results show that FastKV achieves 2.00$\times$ and 1.40$\times$ improvements in time-to-first-token (TTFT) and throughput, respectively, compared to HeadKV, the state-of-the-art KV cache compression method. Moreover, FastKV successfully maintains accuracy on long-context benchmarks at levels comparable to the baselines. Our code is available at https://github.com/dongwonjo/FastKV.
1 Introduction
Figure 1: Comparison of accuracy, TTFT, throughput across different KV cache compression methods on LLaMA-3.1-8B-Instruct
Figure 2: Comparison of the number of tokens processed in each layer/head of LLMs during prefill computation and KV caching across different KV cache compression techniques. As each token produces its corresponding KV, the KV cache size is directly proportional to the number of tokens processed. The blue background box indicates the set of sharing selected token indices.
Recent advancements in large language models (LLMs) have enabled the processing of long-context sequences, such as those comprising 128k tokens Achiam et al. (2023); Gemini Team et al. (2023); Anthropic . This capability significantly broadens the range of applications for LLMs Yi et al. (2024); Laban et al. (2023); Gu (2023). However, as the length of the input sequence increases, the size of these key-value (KV) caches also increases, making them a significant bottleneck in the serving of long-context LLMs. Therefore, compression of KV caches is essential for optimizing the operation of LLMs.
There has been active research on compressing KV cache to alleviate its burden in long-context handling Zhang et al. (2023); Xiao et al. (2023; 2024); Oren et al. (2024); Chen et al. (2024). Some techniques have demonstrated the ability to preserve accuracy comparable to full-context processing after the compression Li et al. (2024); Feng et al. (2024); Fu et al. (2024), while there exists a compression technique that has achieved inference speedup in both the prefill and generation stages Shi et al. (2024). However, no KV cache compression techniques currently exist that can simultaneously preserve accuracy and achieve inference speedup, particularly during the prefill stage.
In this paper, we propose FastKV, an innovative approach designed to expedite long-context handling with LLMs while maintaining their accuracy. FastKV stems from our finding that the properties of attention maps differ between the early and later layers of LLMs. Through detailed analysis of LLMs, we discover that in the later layers, attention maps concentrate on a limited set of significant tokens, which remain consistent across layers. In contrast, the early layers engage with a broader array of tokens and exhibit diverse attention patterns as they process the context of the input prmopt. Based on these findings, FastKV adopts a novel Token-Selective Propagation (TSP) method that applies different strategies of KV cache compression at the early and later layers respectively As shown in Figure 1, this dual-strategy approach enables FastKV to improve both time-to-first-token (TTFT) and throughput of long-context processing while effectively preserves the accuracy of LLMs. Our experimental results show that FastKV can achieve 2.00$\times$ speedup for TTFT and 1.40$\times$ improvement for throughput compared to HeadKV, while maintaining a similar level of accuracy with less than a 1% accuracy gap. These results demonstrate that FastKV promises a practical solution for long context scenarios, particularly in real-time applications that require efficient KV cache management and low latency for the prefill and generation stages.
2 Background
2.1 Long-Context Processing with LLMs
LLMs have significantly enhanced the natural language processing (NLP) abilities of AI systems Brown et al. (2020); Radford (2018); Kamalloo et al. (2023); Jiang et al. (2021) through the use of attention mechanisms Vaswani (2017). These attention mechanisms carefully evaluate the relationships between tokens within sentences to construct context vectors by aggregating token data according to the attention scores to extract context-specific information Brauwers & Frasincar (2021). Distinct from earlier NLP models such as RNNs Sherstinsky (2020) and LSTMs Hochreiter (1997), attention mechanisms consider the entire token data from input sequences, thereby avoiding issues of prompt forgetting. Benefiting from these advancements, recent researches indicates that LLMs are capable of processing long-context sequences.
However, the very nature of LLMs, which retain all token data to facilitate attention mechanisms, leads to significant memory overhead. Typically, token data for attention mechanisms are stored in the form of KV cache, and the size of KV cache escalates with sequence length, becoming a primary source of inefficiency in LLMs when processing long-context scenarios. For instance, in the LLaMA-3.1-8B model Dubey et al. (2024), managing a sequence of 128k tokens requires 17.12GB of memory for KV cache, whereas the model itself only requires 16.06GB. Therefore, the excessive KV cache demands form a critical bottleneck in LLM inference during long-context processing, and efficient compression of KV cache is crucial for streamlined LLM operation.
Algorithm 1 FastKV algorithm for the KV cache compression during the prefill stage
1:$input$ $sequence$$\{I\},$$\#layers$$\{L\},$ $TSPlayer,$ $TSPlength$$\{B_{TSP}\},$ $KVbudget$$\{B_{KV}\}$
2:$generated$ $token$$\{O\},$ $KV$ $Cache$$\{C\}$
3:$X\leftarrow Embedding(I)$
4:for $l$ = 0 to $L-1$ do
5:     if $l\leq TSPlayer$ then
6:         $X,Att_{l},K_{X},V_{X}\leftarrow layer_{l}(X)$
7:         $K,V\leftarrow KVCompress(K_{X},V_{X},Att_{l},B_{KV})$
8:         if $l==TSPlayer$ then
9:              $x\leftarrow HiddenCompress(X,Att_{l},B_{TSP})$
10:         end if
11:     else
12:         $x,Att_{l},K_{x},V_{x}=layer_{l}(x)$
13:         $K,V\leftarrow KVCompress(K_{x},V_{x},Att_{l},B_{KV})$
14:     end if
15:     $C\leftarrow update(K,V)$
16:end for
17:$O\leftarrow LMHead(x)$
18:return $O,C$
2.2 KV Cache Compression
KV cache compression techniques for long-context processing are primarily designed to prune tokens, thereby reducing the associated KV caches and effectively managing memory usage. A common approach leverages the attention scores generated by each attention head to selectively prune tokens in an attention-head-wise manner. This method operates on the premise that token importance varies across attention heads. A notable example is SnapKV Li et al. (2024) in Figure 2(a). It discovers that each attention head in the model consistently focuses on specific prompt token during generation, and this robust pattern can be obtained from an observation window located at the end of the input prompt. Thus, SnapKV evaluates token importance score for each attention head by summing the attention scores assigned to each token as follows:
$S^{l,h}_{i}=\frac{1}{2w_{p}+1}\sum_{m=-w_{p}}^{w_{p}}\sum_{n=0}^{N_{\text{obs}}}Att_{l}[h,\,N_{I}-n,\,i+m]$ (1)
Here, $S_{i}^{l,h}$ is the importance score of $i$-th token in $h$-th attention head of the $l$-th layer. $Att_{l}$ denotes the attention score matrix of $l$-th layer, while $N_{I}$ and $N_{obs}$ indicate the number of tokens in the input prompt and the observation window, respectively. SnapKV applies average pooling on the scores to prevent sparse token selection. Here, $12w_{p}+$ 1 indicates the size of pooling window. SnapKV selects tokens from the observation window first and then selects tokens with the highest importance scores for KV caching. Since SnapKV relies on the attention scores produced by each attention head, it requires propagationg of the full input prompt throughout the entire model, during the prefill stage. As a result, SnapKV cannot improve computational efficiency or reduce the latency of the prefill stage, which directly impacts TTFT, a key factor in optimizing user experience for LLM serving.
AdaKV Feng et al. (2024) in Figure 2(b) expands SnapKV approach to grouped-query-attention (GQA) mechanism Ainslie et al. (2023) by introducing dynamic KV budget allocation across attention groups and selecting important tokens in attention-group-wise manner. To enhance accuracy preservation, HeadKV Fu et al. (2024) in Figure 2(c) builds upon SnapKV by introducing dynamic KV budget allocation across attention heads. This feature has shown effective in preserving accuracy under aggressive KV cache compression scenarios, such as when the average KV budget is as low as 128. However, when applied to more realistic KV budget targets, such as 512, dynamic KV budget allocation exhibits marginal accuracy improvements or even a slight accuracy degradation compared to SnapKV, while they introduces notable latency and throughput overhead. The primary reason of the overhead is that the processing speed of the attention mechanism is limited by the attention head with the longest KV cache. As the size of KV budget increases, the imbalance in KV cache sizes across attention heads becomes more pronounced and it amplifies latency overhead. This overhead makes these methods less practical for real-world applications, where achieving low latency is critical.
In contrast, GemFilter Shi et al. (2024) in Figure 2(d) adopts a fundamentally different approach that enhances the computational efficiency of the prefill stage. GemFilter is based on the observation that LLMs can identify relevant tokens in intermediate layers, referred to as the filter layer. The filter layer selects the indices of relevant tokens at the layer level rather then head level, and GemFilter uses these indices to compress the input prompt. Then, the prefill stage is recomputed with compressed input prompt. By processing full-context with part of the LLM, GemFilter effectively reduces TTFT. The GemFilter paper reports successful accuracy preservation when using a KV cache budget of 2k, where the input prompt is compressed with 2k tokens. Moreover, as GemFilter focuses on the selected tokens more intensively than a full-context model by only processing selected tokens, with sufficient KV budget, there are some cases where it achieves even better accuracy than the full-context model. However, as the removal of tokens from the input prompt results in the complete loss of any information embedded in the discarded tokens, GemFilter struggles to maintain accuracy with lower KV budgets, such as 512, and even with 2k KV budgets when handling complex tasks. Thus, there are currently no KV cache compression techniques that can adequately preserve accuracy while simultaneously improving TTFT/throughput. This underscores the need for new methods that effectively balance these objectives.
3 Proposed FastKV
3.1 Overview of FastKV
The overview of FastKV is illustrated in Figure 2(e) and Algorithm 1. FastKV implements distinct KV cache compression strategies for the early and later layers of LLMs. To speedup the prefill stage, FastKV employs the TSP approach in the later layers. This approach is based on the insight that after the early layers have analyzed contextual information, the later layers tend to focus consistently on the same crucial tokens. The TSP approach, therefore, involves selectively propagating only a limited set of tokens (e.g. 2k tokens) identified by the TSP layer, which is strategically positioned in the middle of the LLM architecture. The TSP layer utilizes its attention map to pinpoint these critical tokens. Unlike the early layers, which propagate the entire set of tokens, the TSP approach in the later layers transmits only the selected tokens to subsequent layers. This targeted propagation means that the later layers process far fewer tokens, thus it can accelerate the processing speed in these layers. While the TSP approach results in a smaller set of KV data for the selected tokens, FastKV still applies the same KV cache compression techniques used in the early layers to optimize the compression ratio. In the early layers, FastKV evaluates the attention scores generated by each attention head to determine the importance of KV data. GQA compatible KV cache compression of FastKV selectively removes the KV data of tokens that receive low attention scores in an attention-group-wise manner. It is important to note that KV data compression occurs solely during the saving of KV data to establish KV caches for input prompt. Consequently, FastKV ensures that the entire set of KV data, including those of the removed tokens, is propagated to the subsequent layer until the TSP layer. This approach guarantees full context propagation to preserve essential contextual information from the input prompt. This dual strategy enables FastKV to achieve significant speed improvements without compromising accuracy, as it properly preserves the contextual information of input sequences.
3.2 Proposed TSP
TSP aims to enhance the efficiency of the prefill stage by propagating only a selected set of important tokens, similar to the approach used by GemFilter. However, in response to the insights obtained from the previous sections, TSP is designed to retain the necessary scope of token propagation to ensure that critical contextual information is preserved. The operation of TSP is governed by three key factors: the layer for token selection (TSP layer), identification of important tokens, and the number of tokens selected for propagation (TSP length). TSP begins by identifying TSP length important tokens from the full tokens at the TSP layer. The method for identifying important tokens in TSP builds upon Equation 1. The propagation of selected tokens across LLM layers involves transmitting the hidden states associated with those selected tokens. Consequently, TSP necessitates the assessment of token importance on a layer-by-layer basis. To facilitate this, TSP includes an additional step that aggregates the attention scores from all attention heads within the TSP layer as follows:
$S_{i}^{TSPlayer}=\frac{1}{H}\sum_{h=0}^{H-1}S_{i}^{TSPlayer,h}$ (2)
Here, $H$ denotes the number of attention head in each layer. From the TSP layer onward, TSP propagates only selected important tokens through the later layers. Unlike GemFilter, which reverts to the first layer of the LLM with selected tokens and recomputes the prefill stage, TSP directly propagates the selected tokens to subsequent layers without returning to the first layer. Meanwhile, although previous approaches to KV cache compression aimed at high accuracy by propagating the entire context throughout the prefill stage, the proposed TSP method discards the unimportant tokens after the TSP layer. Therefore, the implementation of TSP must be carefully designed based on a detailed analysis of the properties of token importance and its impact on the LLM output.
3.3 Impact of Important Tokens
Figure 3: Rate of sum of attention scores for the top 2k important tokens selected to the total attention score at each layer.
In the LLM, the attention mechanism is the critical component for extracting contextual information, while other parts of the model process input in a more token-isolated manner. Within the attention mechanism, each token generates a new embedding by integrating the embeddings of all preceding tokens, with this integration weighted by the attention scores. Consequently, tokens that receive low attention scores have only a marginal impact on the processing of contextual information. To ensure the removal of the unimportant tokens without significant impact on the LLM output, the important tokens must receive significantly higher attention scores.
We evaluate the actual impact of these important tokens by examining their attention scores relative to those of the remaining tokens, as shown in Figure 3. In this analysis, we calculate the sum of the attention scores assigned to 2k important tokens selected in each layer and compare this sum to the attention scores allocated to all 128k tokens. Despite important tokens comprising only 1.56% of the total number of tokens, they receive a majority of the attention scores. This is particularly notable in the later layers, where the proportion of attention scores attributed to these tokens consistently exceeds 50% and often approaches 80%. This result indicates that they are central to the LLM processing and the remaining tokens could be removed with a marginal impact on the LLM output.
3.4 Dynamics of Token Importance
Figure 4: Minimum match rate between indices of 2k important tokens selected at each layer and its subsequent layers.
Figure 5: Visualization of output logits by t-SNE for (left) LLaMA-3.1-8B-Instruct and (right) Mistral-Nemo-12B-Instruct.
Given that tokens discarded after the TSP layer are not propagated in subsequent layers, there must be consistency in identifying important tokens across layers to facilitate TSP. This section examines the dynamics of token importance across layers of LLMs. Specifically, we investigate whether tokens deemed important based on their high attention scores remain consistent across LLM layers. A significant overlap in the set of important tokens would suggest the potential for pre-selecting these tokens before propagating them to subsequent layers. To evaluate the consistency of important tokens across layers, we analyze the important tokens in each layer of the LLM and calculate the minimum match rate of the indices of these tokens with those in subsequent layers. The minimum match rate serves as an indicator of an information bottleneck.
Figure 4 presents our findings from this analysis on various LLMs, where each layer extracts 2k important tokens from 128k tokens. The minimum match rate in the first layer is below 10%, indicating low correlation between important tokens in the first and later layers. However, as we progress to subsequent layers, this match rate increases, surpassing 25% by the middle layer. This indicates that among the 2k tokens, at least 512 consistently appear as important across layers. Therefore, if the KV cache budget per layer is set to 512, it suffices to examine the 2k important tokens selected from the middle layer. The observed consistency in important token selection in later layers likely stems from the increasing robustness of contextual information extraction. Each layer in an LLM uses attention scores to integrate token information, progressively enhancing contextual understanding. As the model advances through its layers, contextual analysis becomes more refined, enabling a more precise identification of relevant tokens. Thus, later layers of LLMs exhibit a consistent recognition of important tokens, underscoring their refined understanding of token importance in the context of the overall sequence.
3.5 Overall Impact of TSP on LLM Output
Figure 6: LongBench results of LLaMA-3.1-8B-insturct with propagation of tokens selected at each layer.
Building on the insights from Section 3.3 and 3.4, we can expect that the removal of unimportant tokens identified from the middle layer will not significantly impact the performance of the LLM for two main reasons. First, only a small set of important tokens plays an important role in the attention mechanism and there is the consistency in the identification of important tokens in the middle layer and subsequent layers. We analyze real changes in the output logits of an LLM after selecting 2k tokens from each layer and discarding the unselected tokens before propagating them to subsequent layers. For this analysis, we use t-SNE (t-distributed Stochastic Neighbor Embedding) to project the final output logits of LLMs into a 2D space. Figure 5 demonstrates the distances between output logits obtained with full-context propagation and those obtained through propagation focusing exclusively on important tokens. This visual representation shows that removing tokens in the early layers causes significant deviations in output logits compared to full-context propagation. However, as expected, removing tokens from later layers preserves a high degree of similarity to the results from full-context models. This contributes to stable accuracy results on the LongBench benchmark, as illustrated in Figure 6. In light of this observation, we designate the middle layer of each LLM as the TSP layer for the remainder of this paper.
Moreover, in Figure 6, we compared the accuracy obtained with TSP and a GemFilter-like propagation. TSP consistently outperforms GemFilter in terms of accuracy. This superior performance can primarily be attributed to how TSP manages the variability in token importance across layers. Unlike GemFilter, which reverts to the first layer of the LLM with selected tokens and recomputes the prefill stage, TSP directly propagates the selected tokens to subsequent layers without reverting to the initial layer. This direct propagation allows TSP to better accommodate the variability in token importance. Furthermore, the large accuracy gap between TSP and GemFilter when tokens are selected in the initial layers highlights the importance of leveraging token information that has been processed with the full context before making important token selections.
3.6 TSP Length and KV Budget
The match rate of the important tokens is around 25% (Figure 4). This rate is sufficiently high to justify the use of TSP approach with an adequate TSP length (e.g. 2k). However, a short TSP length (e.g. 512) may lead to significant information loss. Therefore, TSP length, which defines the number of tokens selected for propagation, is strategically designed to operate independently of the KV cache budget. For instance, if the KV cache budget is set to 512 while the TSP length is 2048, TSP propagates 2048 tokens during its operation, but only the final 512 tokens are stored in the KV cache. This separation allows TSP to process a significantly larger set of tokens during propagation, enabling it to capture more contextual information while maintaining the KV cache within its predefined budget. As shown in Figure 7, a TSP length of 2048 has been found sufficient to preserve accuracy of the model under specific KV budget while maintaining efficiency in terms of TTFT.
Table 1: LongBench results comparison between baseline methods and FastKV.
Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RB-P
Avg.
LLaMA-3.1-8B-Inustruct, KV Budget = Full
Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 48.63
LLaMA-3.1-8B-Instruct, KV Budget = 512
SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 46.60
AdaKV 29.07 40.16 52.44 53.90 43.05 31.10 25.75 24.39 24.85 69.00 92.34 42.05 63.43 55.32 46.20
HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 47.76
GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 35.64
FastKV 30.38 41.12 55.71 54.35 46.69 30.89 26.78 23.82 24.37 72.50 92.04 42.64 62.60 52.74 46.90
LLaMA-3.1-8B-Instruct, KV Budget = 2048
SnapKV 31.31 44.96 55.15 55.48 45.29 30.78 31.92 24.63 26.97 71.50 91.48 43.38 63.14 56.04 48.00
AdaKV 30.64 45.23 55.46 55.55 45.26 31.14 31.28 24.89 26.72 72.50 91.64 42.75 64.88 59.25 48.37
HeadKV 30.53 44.99 54.98 55.47 46.08 31.42 32.41 25.05 27.20 72.50 91.57 42.84 63.45 56.75 48.23
GemFilter 23.64 41.17 51.39 53.97 45.32 29.24 32.01 20.50 26.91 70.00 91.59 42.59 47.35 38.91 43.90
FastKV 30.31 45.41 54.79 55.11 46.57 30.46 31.25 24.82 27.02 73.50 91.48 43.66 63.17 55.82 48.10
Mistral-Nemo-12B-Instruct, KV Budget = Full
Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 48.33
Mistral-Nemo-12B-Instruct, KV Budget = 512
SnapKV 23.58 40.12 55.17 47.91 45.69 25.17 23.77 22.48 24.08 74.00 89.44 43.09 68.31 62.15 46.07
AdaKV 24.46 41.52 57.34 47.89 45.83 25.57 23.58 22.11 23.88 74.00 89.52 43.20 68.21 64.15 46.52
HeadKV 25.88 40.57 57.12 48.11 46.11 25.26 26.19 22.92 25.25 74.50 89.74 41.81 68.79 64.76 46.93
GemFilter 25.45 37.64 53.91 52.83 51.26 33.02 26.23 19.42 23.87 65.50 84.16 40.05 38.03 41.13 42.32
FastKV 25.51 41.48 56.76 48.73 45.43 26.53 23.42 22.15 24.02 74.00 89.42 43.22 68.61 63.57 46.63
Mistral-Nemo-12B-Instruct, KV Budget = 2048
SnapKV 24.35 42.95 56.92 49.08 46.43 26.26 28.27 23.68 26.07 75.00 89.82 44.20 68.89 68.00 47.85
AdaKV 26.02 43.08 57.43 48.76 46.53 26.12 27.80 23.44 25.93 75.00 89.66 44.49 68.53 67.57 47.88
HeadKV 26.05 42.15 57.46 48.49 45.85 25.86 30.27 24.22 26.17 75.00 89.66 43.74 68.64 67.29 47.92
GemFilter 26.42 42.40 56.98 57.64 53.92 34.33 30.51 21.60 25.96 72.00 89.65 44.48 48.34 48.06 46.59
FastKV 26.61 43.37 57.12 49.41 47.22 26.27 28.40 23.22 25.81 75.00 89.72 43.04 68.67 67.61 48.11
3.7 GQA Compatible KV Cache Compression
While FastKV incorporates TSP for the latter half of the LLM, it takes a different approach to KV cache compression by propagating full-context data in the early layers and evaluating the importance of KV data using attention scores generated by each layer. As the state-of-the-art LLMs Yang et al. (2024); Abdin et al. (2024) increasingly adopt GQA Ainslie et al. (2023), which groups multiple attention heads together and shares KV data within each group, FastKV incorporates GQA compatible KV cache compression technique presented in AdaKV to the early layers. This technique adapts SnapKV approach (Equation 1) to GQA by aggregating attention scores across all heads within the group for each token as follows:
$S_{i}^{l,g}=\frac{1}{H_{G}}\sum_{h=h_{g}}^{h_{g}+H_{G}}S_{i}^{j,h}$ (3)
Here, $H_{G}$ denotes the number of attention heads in each group. $g$ and $h_{g}$ denotes the group index and the index of the first head in the group. This aggregation evaluates the importance of tokens at the group level. FastKV then selects and stores KV cache entries corresponding to the top-KV budget important tokens in an attention-group-wise manner.
Figure 7: LongBench results (line) and TTFT (bar) of LLaMA-3.1-8B-insturct across various KV budgets and TSP lengths.
Figure 8: Needle-in-a-Haystack results of LLaMA-3.1-8B-Instruct with 512 KV budget.
4 Experiments
4.1 Setup
Models & Datasets. We evaluate two open-source LLMs of different sizes: LLaMA-3.1-8B-Instruct Dubey et al. (2024) and Mistral-Nemo-12B-Instruct Mistral AI Team (2024). These models have 32 and 40 decoder layers, respectively, and employ GQA Ainslie et al. (2023) with a context window size of 128k tokens.
To assess long-context understanding and retrieval capabilities, we use two benchmark datasets: LongBench Bai et al. (2023) and Needle-in-a-Haystack Kamradt (2023).
Implementation Details. We integrated our proposed FastKV method upon self-attention implementation of HuggingFace Transformers library, which utilizes FlashAttention-2 Dao (2023) kernel. We select layer 15 and 19 as the TSP layer for LLaMA-3.1-8B-Instruct, and Mistral-Nemo-12B-Instruct, respectively. The TSP length is set to 2048. We fix the observation window size ($N_{obs}$) to 8 and the kernel size of average pooling ($12w_{p}+$ 1) to 7.
Baselines. We compare our approach with four baseline methods for KV cache compression: SnapKV Li et al. (2024), AdaKV Feng et al. (2024), HeadKV Fu et al. (2024), and GemFilter Shi et al. (2024). SnapKV, AdaKV, and HeadKV use the same local window size, pooling method, and pooling kernel size as FastKV. The indices of GemFilter filter layers for LLaMA-3.1-8B-Instruct and Mistral-Nemo-12B-Instruct are 13 and 19, respectively Shi et al. (2024).
4.2 Accuracy Evaluation
LongBench. The accuracy evaluation results on LongBench are summarized in Table 1. The full breakdown of results is provided in Appendix A.1. Previous works, such as SnapKV, AdaKV, and HeadKV, which use attention-head-wise or attention-group-wise token removal, successfully maintain accuracy after KV cache compression, with an average accuracy drop of less than 2.5% and 0.51% compared to FullKV for 512 and 2048 KV budgets, respectively. GemFilter, which discards unselected tokens from the input prompt and recomputes the prefill stage to achieve both memory and computational efficiency, results in a significantly larger accuracy drop, with reductions of up to 13.0% and 4.7% compared to FullKV for 512 and 2048 KV budgets, respectively. On the other hand, while the proposed FastKV is designed to achieve memory and computational efficiency comparable to GemFilter, it successfully preserves accuracy on par with SnapKV, AdaKV, and HeadKV. It presents the effectiveness of FastKV’s distinct KV cache compression strategies for the early and later layers of LLMs.
Needle-in-a-Haystack. Figure 8 presents the Needle-in-a-Haystack evaluations for LLaMA-3.1-8B-Instruct with 512 KV budget. Additional analysis results with various LLM models and KV budgets can be found in Appendix A.2. Consistent with the results observed in LongBench, GemFilter shows the worst performance, while FastKV achieves the best performance with a slight improvement over HeadKV.
Figure 9: (a) TTFT and (b) throughput results across different methods on LLaMA-3.1-8B-Instruct. (dashed line: Full KV)
4.3 Latency and Throughput Evaluation
We evaluate the latency of the prefill stage (time-to-first-token (TTFT)) and throughput of token generation with LLaMA-3.1-8B-Instruct model and a single NVIDIA A100 GPU. Figure 9 reports TTFT results with 64k/128k input with 512 KV budget, and throughput results with 128k input with 512/2048 KV budget. More detailed evaluation results can be found in Appendix A.3.
TTFT. As shown in Figure 9(a), both GemFilter and FastKV achieve significant TTFT improvements over other baselines by reducing the computational complexity of the prefill stage through the concept of propagating only selected tokens during this stage. As a result, FastKV achieves TTFT that is 1.60$\times$ shorter and 1.97$\times$ shorter than Full KV, with 64k and 128k input, respecitvely. While both methods enhance efficiency, GemFilter incurs additional overhead due to the need for prefill recomputation, whereas FastKV introduces extra processes for compressing generated KV caches. These two factors offset each other. Consequently, GemFilter’s slight advantage in TTFT is primarily attributed to its earlier filter layer (13) compared to the TSP layer (15) for LLaMA-3.1-8B-Instruct. In contrast, SnapKV, AdaKV, and HeadKV exhibit even longer TTFT than Full KV. This is because they process the full-context information during the prefill stage, similar to Full KV, while also incurring additional overhead from indexing and selecting crucial tokens.
Throughput. We measure the throughput for generating a total of 128 tokens. As shown in Figure 9(b), all KV compression techniques result in throughput improvements compared to Full KV. Both FastKV and GemFilter achieve a significant level of throughput improvement. FastKV’s throughput reaches 5.07$\times$ of Full KV and 1.40$\times$ of HeadKV with 512 KV budget. In contrast, SnapKV, which stores KV caches of different tokens set for each attention head, shows slightly lower throughput. Additionally, HeadKV, which allocates varying KV budgets across attention heads, exhibits the lowest throughput among KV cache compression techniques. Managing different sequence lengths for keys and values across attention heads increases the complexity of the attention mechanism. This effect gets stronger when the KV cache size is large, thereby HeadKV experiences steep decline in throughput as KV budget increases, whereas other methods’ throughput remain relatively consistent. AdaKV also adopts fine-grained KV budget allocation similar to HeadKV except that its granularity is attention-group. AdaKV mitigates the influence of fragmented KV budget to some extent thanks to GQA-aware compression, but it does not achieve the same level of throughput with FastKV or GemFilter.
5 Conclusion
In this paper, we introduce FastKV, a novel KV cache compression method designed to improve the efficiency of long-context LLM inference while preserving accuracy. Unlike previous KV cache compression techniques, FastKV enhances computational efficiency by introducing Token-Selective Propagation (TSP). This method strategically retains full-context information in early layers and propagates a compressed subset of tokens in later layers, reducing the computational cost of the prefill stage. Furthermore, our approach is optimized by integrating GQA compatible KV cache compression, which enables more efficient memory and computation management by leveraging group-wise token selection. Our experimental results demonstrate that FastKV achieves higher throughput and lower TTFT than baseline methods while maintaining high accuracy on long-context benchmarks.
Appendix A More Experimental Results
A.1 LongBench
We provide a full breakdown of the LongBench evaluation results for LLaMA-3.1-8B-Instruct (Table 2), LLaMA-3.2-3B-Instruct (Table 3), and Mistral-Nemo-12B-Instruct (Table 4). For LLaMA-3.2-3B-Instruct, which has 28 decoder layers, we set the TSP layer and GemFilter filter layer to 13 both.
Table 2: LongBench results of LLaMA-3.1-8B-Instruct.
Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RB-P
Avg.
LLaMA-3.1-8B-Instruct, KV Cache Size = Full
Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 48.63
LLaMA-3.1-8B-Inustruct, KV Budget = 128
SnapKV 27.39 30.40 50.35 53.04 43.49 28.89 22.87 23.00 21.74 62.00 91.19 40.31 58.43 49.33 43.03
AdaKV 24.90 24.41 49.95 53.15 41.73 28.55 20.54 23.21 20.28 50.50 89.49 40.71 58.74 52.40 41.33
HeadKV 29.74 40.34 53.51 54.83 45.35 30.62 26.40 24.23 24.58 71.50 88.51 40.14 60.19 52.92 45.92
GemFilter 12.50 11.00 16.52 28.84 18.82 13.53 19.66 17.67 14.34 58.17 78.51 32.99 19.92 21.35 25.99
FastKV 27.63 27.84 52.32 52.00 45.27 28.35 22.00 22.81 20.64 69.50 91.28 40.17 59.08 48.38 42.66
LLaMA-3.1-8B-Instruct, KV Budget = 512
SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 46.60
AdaKV 29.07 40.16 52.44 53.90 43.05 31.10 25.75 24.39 24.85 69.00 92.34 42.05 63.43 55.32 46.20
HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 47.76
GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 35.64
FastKV 30.38 41.12 55.71 54.35 46.69 30.89 26.78 23.82 24.37 72.50 92.04 42.64 62.60 52.74 46.90
LLaMA-3.1-8B-Instruct, KV Budget = 1024
SnapKV 31.23 42.52 53.96 55.48 45.43 31.50 29.54 24.78 26.17 70.50 91.73 42.52 62.25 54.72 47.31
AdaKV 29.23 44.09 53.82 54.80 44.01 31.40 28.86 24.73 26.04 72.50 91.72 42.56 63.22 56.33 47.38
HeadKV 30.55 44.66 54.69 55.47 46.20 31.63 32.88 25.14 27.10 73.00 91.57 42.88 63.76 56.06 48.26
GemFilter 19.34 34.76 46.58 46.82 42.83 27.51 29.89 18.96 25.68 63.00 90.70 42.50 38.09 35.14 40.13
FastKV 30.22 44.01 55.04 54.76 46.76 30.59 29.43 24.66 26.19 73.50 91.76 42.42 62.54 55.02 47.64
LLaMA-3.1-8B-Instruct, KV Budget = 2048
SnapKV 31.31 44.96 55.15 55.48 45.29 30.78 31.92 24.63 26.97 71.50 91.48 43.38 63.14 56.04 48.00
AdaKV 30.64 45.23 55.46 55.55 45.26 31.14 31.28 24.89 26.72 72.50 91.64 42.75 64.88 59.25 48.37
HeadKV 30.53 44.99 54.98 55.47 46.08 31.42 32.41 25.05 27.20 72.50 91.57 42.84 63.45 56.75 48.23
GemFilter 23.64 41.17 51.39 53.97 45.32 29.24 32.01 20.50 26.91 70.00 91.59 42.59 47.35 38.91 43.90
FastKV 30.31 45.41 54.79 55.11 46.57 30.46 31.25 24.82 27.02 73.50 91.48 43.66 63.17 55.82 48.10
Table 3: LongBench results of LLaMA-3.2-3B-Instruct.
Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RB-P
Avg.
LLaMA-3.2-3B-Instruct, KV Budget = Full
Full KV 23.41 40.60 49.79 48.71 38.89 20.54 34.07 23.65 26.67 71.50 88.89 43.20 52.13 54.16 44.02
LLaMA-3.2-3B-Instruct, KV Budget= 128
SnapKV 17.92 23.18 44.59 45.35 35.08 17.07 21.96 21.05 20.38 66.00 87.55 38.39 48.31 45.82 38.05
AdaKV 19.94 23.15 44.50 45.17 36.51 18.68 21.86 21.46 20.57 67.00 87.82 39.26 49.39 47.61 38.78
HeadKV 21.59 32.15 48.00 48.79 38.28 18.51 25.30 21.94 23.86 69.50 88.45 38.34 52.11 50.25 41.22
GemFilter 10.36 10.70 26.36 35.84 27.75 11.63 19.23 17.00 13.59 55.50 71.58 28.40 19.13 27.46 26.75
FastKV 19.95 23.09 44.34 45.25 34.49 19.89 21.34 20.63 20.02 66.00 88.05 38.37 48.91 47.13 38.39
LLaMA-3.2-3B-Instruct, KV Budget = 512
SnapKV 21.03 32.79 49.01 48.68 37.59 20.07 25.91 22.76 24.22 69.00 88.75 41.78 51.68 51.25 41.75
AdaKV 21.30 32.88 49.92 47.65 37.20 19.72 25.96 22.65 24.31 69.50 89.03 41.55 51.63 51.87 41.80
HeadKV 22.79 36.45 50.57 48.85 39.16 20.44 29.79 23.24 26.10 70.50 89.66 40.80 53.72 53.83 43.28
GemFilter 20.88 31.37 44.31 47.60 34.87 23.59 25.89 19.82 24.01 55.00 85.73 35.29 26.68 32.72 36.27
FastKV 21.73 31.55 49.88 47.31 36.54 22.14 25.47 22.38 24.05 71.50 88.99 40.91 52.46 52.17 41.93
LLaMA-3.2-3B-Instruct, KV Budget = 1024
SnapKV 21.71 36.22 49.04 48.78 38.14 21.21 28.32 23.25 25.62 69.00 89.03 41.68 53.01 53.42 42.75
AdaKV 21.69 36.91 49.76 48.10 39.27 21.25 28.01 22.78 25.66 70.00 88.96 42.62 53.00 53.35 42.96
HeadKV 23.44 38.06 50.37 49.14 39.51 20.33 31.48 23.26 26.59 71.00 89.53 41.41 53.84 54.14 43.72
GemFilter 21.34 33.72 48.52 50.13 33.75 20.82 28.90 20.73 24.73 61.50 89.17 38.44 29.89 34.05 38.26
FastKV 23.13 35.27 49.48 47.93 38.74 21.42 28.09 22.66 25.86 71.50 89.46 41.46 53.09 54.37 43.03
LLaMA-3.2-3B-Instruct, KV Budget = 2048
SnapKV 22.89 38.70 50.91 48.38 39.29 20.55 30.79 23.17 26.46 70.50 88.89 42.25 52.16 54.03 43.50
AdaKV 23.43 39.28 50.26 48.61 39.60 20.82 30.39 22.94 26.65 71.00 88.89 42.57 52.33 54.60 43.67
HeadKV 22.56 39.84 50.40 49.15 38.95 20.25 32.98 23.73 26.71 71.00 89.28 42.74 53.56 55.30 44.03
GemFilter 17.48 38.31 49.93 50.57 40.29 19.02 30.95 21.50 26.28 64.50 90.07 40.59 37.97 38.88 40.45
FastKV 24.02 37.99 50.06 47.58 38.76 21.93 30.45 23.24 26.29 72.50 89.22 41.73 52.71 55.52 43.71
Table 4: LongBench results of Mistral-Nemo-12B-Instruct.
Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RB-P
Avg.
Mistral-Nemo-12B-Instruct, Budget = Full
Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 48.33
Mistral-Nemo-12B-Instruct, KV Budget = 128
SnapKV 21.29 34.24 51.53 48.07 44.55 23.99 19.93 20.73 20.69 69.50 89.52 39.92 62.85 53.08 42.85
AdaKV 22.29 36.52 52.50 47.28 44.73 21.84 19.15 20.61 20.74 65.50 88.64 39.25 63.79 54.07 42.64
HeadKV 22.62 38.76 52.54 46.49 45.97 21.80 20.74 21.37 22.10 73.00 89.42 39.16 64.92 56.15 43.93
GemFilter 15.74 18.30 42.98 41.86 34.20 18.44 21.72 16.56 18.81 57.00 63.02 31.78 25.74 30.62 31.20
FastKV 22.26 35.09 53.33 49.15 45.63 23.67 19.15 20.80 20.47 65.00 90.06 39.88 63.65 55.68 43.13
Mistral-Nemo-12B-Instruct, KV Budget = 512
SnapKV 23.58 40.12 55.17 47.91 45.69 25.17 23.77 22.48 24.08 74.00 89.44 43.09 68.31 62.15 46.07
AdaKV 24.46 41.52 57.34 47.89 45.83 25.57 23.58 22.11 23.88 74.00 89.52 43.20 68.21 64.15 46.52
HeadKV 25.88 40.57 57.12 48.11 46.11 25.26 26.19 22.92 25.25 74.50 89.74 41.81 68.79 64.76 46.93
GemFilter 25.45 37.64 53.91 52.83 51.26 33.02 26.23 19.42 23.87 65.50 84.16 40.05 38.03 41.13 42.32
FastKV 25.51 41.48 56.76 48.73 45.43 26.53 23.42 22.15 24.02 74.00 89.42 43.22 68.61 63.57 46.63
Mistral-Nemo-12B-Instruct, KV Budget = 1024
SnapKV 23.27 41.42 56.01 47.85 46.09 25.83 25.97 23.70 25.37 74.00 89.82 43.48 68.46 65.84 46.94
AdaKV 25.33 41.82 57.52 48.02 46.48 26.34 25.68 23.12 25.26 74.50 89.82 44.65 68.49 66.21 47.37
HeadKV 25.94 41.38 56.45 48.22 45.90 25.05 28.67 23.72 25.72 75.00 89.84 42.64 68.66 66.60 47.41
GemFilter 26.94 40.67 53.63 55.82 56.90 38.23 28.68 20.76 25.34 69.00 87.32 42.49 44.33 45.41 45.39
FastKV 26.26 42.37 57.71 49.40 46.88 25.58 26.17 23.08 25.29 75.00 89.72 43.56 68.53 65.79 47.52
Mistral-Nemo-12B-Instruct, KV Budget = 2048
SnapKV 24.35 42.95 56.92 49.08 46.43 26.26 28.27 23.68 26.07 75.00 89.82 44.20 68.89 68.00 47.85
AdaKV 26.02 43.08 57.43 48.76 46.53 26.12 27.80 23.44 25.93 75.00 89.66 44.49 68.53 67.57 47.88
HeadKV 26.05 42.15 57.46 48.49 45.85 25.86 30.27 24.22 26.17 75.00 89.66 43.74 68.64 67.29 47.92
GemFilter 26.42 42.40 56.98 57.64 53.92 34.33 30.51 21.60 25.96 72.00 89.65 44.48 48.34 48.06 46.59
FastKV 26.61 43.37 57.12 49.41 47.22 26.27 28.40 23.22 25.81 75.00 89.72 43.04 68.67 67.61 48.11
A.2 Needle-in-a-Haystack
We performed further Needle-in-a-Haystack evaluations with varying KV budgets for LLaMA-3.1-8B-Instruct (Figure 10), LLaMA-3.2-3B-Instruct (Figure 11), and Mistral-Nemo-12B-Instruct (Figure 12). For LLaMA-3.1-8B-Instruct, the overall retrieval trend remains consistent. FastKV continues to achieve strong performance across multiple budgets, often matching or surpassing baseline methods, including HeadKV.
By contrast, LLaMA-3.2-3B-Instruct produces notably lower scores overall, even when using Full KV. Its smaller parameter count appears to limit retrieval capabilities at long contexts, making it more difficult for fine-grained KV allocation techniques to improve performance. Although HeadKV excelled with LLaMA-3.1-8B-Instruct, it does not show similar gains on LLaMA-3.2-3B-Instruct, suggesting that certain highly flexible optimizations may not translate effectively to smaller models. Nonetheless, FastKV outperforms SnapKV, HeadKV, and GemFilter, and achieves an average score on par with AdaKV, indicating that its core compression strategy retains efficacy when applied to a compact model.
Mistral-Nemo-12B-Instruct exhibits a shorter effective context window than the nominal 128k, frequently struggling to retrieve needles located at extreme depths while handling nearer contexts reasonably well. Surprisingly, GemFilter significantly outperform Full KV, and FastKV also records above-Full-KV scores. GemFilter discards a subset of tokens and restarts its prefill stage with the reduced sequence, and FastKV continues prefilling the remaining layers with reduced hidden states after TSP. This may lead temporary enhancement of retrieval by focusing attention on fewer tokens. However, GemFilter’s poor results in LongBench confirm that these gains do not generalize to complex tasks and models.
Figure 10: Needle-in-a-Haystack results of LLaMA-3.1-8B-Instruct.
Figure 11: Needle-in-a-Haystack results of LLaMA-3.2-3B-Instruct.
Figure 12: Needle-in-a-Haystack results of Mistral-Nemo-12B-Instruct.
A.3 Latency and Throughput Evaluation
We present the TTFT and the throughput results of the baselines and FastKV in Table 5 and Table  7. The throughput results of Full KV is depicted in Table  6 for effective comparison. The throughput results of KV cache compression methods are evaluated with input context length 128k. As the input context length increases, the relative TTFT gain of FastKV compared to SnapKV, AdaKV and HeadKV increases. For Mistral-Nemo-12B-instruct, FastKV achieves TTFT similar to GemFilter because the TSP layer index is same with the GemFilter filter layer. The token throughput gain of the baselines and FastKV using fixed KV budget grows rapidly as the throughput of Full KV declines due to the increasing size of KV cache. FastKV consistently achieves the highest throughput across all scenarios.
Table 5: TTFT (sec) comparison of baselines and FastKV.
Context Length
Method 8k 16k 32k 64k 128k
LLaMA-3.1-8B-Instruct
     Full KV 0.64 1.41 3.40 9.05 27.94
     SnapKV 0.67 1.45 3.49 9.22 28.27
     AdaKV 0.75 1.53 3.57 9.31 28.36
     HeadKV 0.70 1.48 3.51 9.25 28.30
     GemFilter 0.44 0.78 1.66 4.15 12.51
     FastKV 0.41 0.80 1.81 4.66 14.15
Mistral-Nemo-12B-Instruct
     Full KV 0.92 2.04 4.85 12.71 37.94
     SnapKV 0.95 2.08 4.90 12.80 38.11
     AdaKV 1.05 2.18 5.01 12.92 38.23
     HeadKV 0.99 2.12 4.95 12.85 38.18
     GemFilter 0.69 1.25 2.65 6.59 19.22
     FastKV 0.59 1.16 2.57 6.53 19.22
Table 6: Throughput (Tokens/sec) with full KV budget at varying context length.
    Context Length 8k 16k 32k 64k 128k
    LLaMA-3.1-8B-Instruct 51.55 43.07 32.29 21.24 12.60
    Mistral-Nemo-12B-Instruct 36.10 30.88 23.67 16.07 9.68
Table 7: Throughput (Tokens/sec) comparison of baselines and FastKV.
        Method         KV Budget
        512         1024         2048
        LLaMA-3.1-8B-Instruct
        SnapKV         48.71         48.62         48.29
        AdaKV         50.57         49.53         49.47
        HeadKV         45.64         39.48         30.78
        GemFilter         63.40         62.32         60.37
        FastKV         63.84         62.97         60.74
        Mistral-Nemo-12B-Instruct
        SnapKV         38.04         38.03         35.44
        AdaKV         28.53         28.34         27.82
        HeadKV         27.20         26.97         24.92
        GemFilter         41.94         41.69         40.17
        FastKV         42.01         41.85         41.49

### v2

@@
-
+FastKV: KV Cache Compression for Fast Long-Context Processing with Token-Selective Propagation
 Abstract
@@
 Previous efforts to compress these KV caches primarily focused on reducing memory demands but were limited in enhancing latency.
-To address this issue, we introduce FastKV, a KV cache compression method designed to enhance latency for long-context sequences.
-To enhance processing speeds while maintaining accuracy, FastKV adopts a novel Token-Selective Propagation (TSP) approach that retains the full context information in the initial layers of LLMs and selectively propagates only a portion of this information in deeper layers even in the prefill stage.
-Additionally, FastKV incorporates grouped-query attention (GQA)-aware KV cache compression to exploit the advantages of GQA in both memory and computational efficiency.
-Our experimental results show that FastKV achieves 2.00$\times$ and 1.40$\times$ improvements in time-to-first-token (TTFT) and throughput, respectively, compared to HeadKV, the state-of-the-art KV cache compression method.
-Moreover, FastKV successfully maintains accuracy on long-context benchmarks at levels comparable to the baselines.
+To address this issue, we introduce FastKV, a KV cache compression method designed to reduce latency for long-context inference.
+FastKV improves processing speed while preserving accuracy by adopting Token-Selective Propagation (TSP).
+This approach preserves full-context information in early layers of LLMs and selectively propagates only a portion of this information in later layers.
+This design enables FastKV to minimize redundant computation without sacrificing contextual fidelity.
+Our experimental results show that FastKV achieves up to 1.97$\times$ and 4.82$\times$ improvements in time-to-first-token (TTFT) and throughput, respectively, compared to baseline without KV cache compression.
+Moreover, FastKV successfully maintains accuracy within 1% of the baseline on long-context benchmarks.
 Our code is available at https://github.com/dongwonjo/FastKV.
 1 Introduction
-Figure 1: Comparison of accuracy, TTFT, throughput across different KV cache compression methods on LLaMA-3.1-8B-Instruct
+Recent advancements in large language models (LLMs) have enabled the processing of long-context sequences, such as those comprising 128K tokens [4, 5, 6, 9, 23].
+This capability significantly broadens the range of applications for LLMs [31, 32, 33].
+However, as the length of the input sequence increases, the size of these key-value (KV) caches also increases, making them a significant bottleneck in the serving of long-context LLMs.
+Therefore, compression of KV caches is essential for optimizing the operation of LLMs.
+There has been active research on compressing KV cache to alleviate the burden in long-context handling [10, 11, 24, 25, 26, 12].
+Some techniques have demonstrated the ability to preserve accuracy comparable to full-context processing after the compression [13, 14, 15], while there exists a compression technique that has achieved inference speedup in both the prefill and generation stages [16].
+However, no existing KV cache compression techniques can simultaneously preserve accuracy and achieve inference speedup, particularly during the prefill stage.
+To bridge this gap, we propose FastKV, a KV cache compression framework that accelerates long-context processing without compromising accuracy.
+FastKV adopts a novel Token-Selective Propagation (TSP) method that applies different strategies of KV cache compression at the early and later layers respectively.
+TSP enables early layers to attend to the full input context while propagating only a small set of critical tokens to later layers.
+As illustrated in Figure 1(a), this dual-strategy enables FastKV to improve both time-to-first-token (TTFT) and throughput of long-context processing while effectively preserving the accuracy of LLMs.
+Our experiments show that FastKV delivers improvements of up to 1.97$\times$ in TTFT and 4.82$\times$ in throughput with less than a 1% accuracy drop compared to baselines.
+These results demonstrate that FastKV promises a practical solution for long-context scenarios, particularly in real-time applications that require efficient KV cache management and low latency for the prefill and generation stages.
+Figure 1: (a) Comparison of LongBench accuracy, TTFT, throughput across different KV cache compression methods on LLaMA-3.1-8B-UltraLong-1M-Instruct.
+(b) End-to-end latency breakdown measured on a 128K input tokens with 512-token generation.
+SnapKV [13] represents typical previous works that compress KV cache for the generation stage only.
+2 Background
+2.1 Long-Context Processing with LLMs
+LLMs have significantly enhanced the natural language processing (NLP) abilities of AI systems [27, 28, 29, 30] through the use of attention mechanisms [21].
+However, the very nature of LLMs, which retain all token data to facilitate attention mechanisms, leads to significant memory overhead.
+Typically, token data for attention mechanisms are stored in the form of KV cache, and the size of KV cache escalates with sequence length, becoming a primary source of inefficiency in LLMs when processing long-context scenarios.
+For instance, in the LLaMA-3.1-8B model [7], managing a sequence of 128K tokens requires 17.18GB of memory for KV cache, whereas the model itself only requires 16.06GB.
+These excessive KV cache demands form a critical bottleneck in LLM inference during long-context processing, and efficient compression of KV cache is crucial for streamlined LLM operation.
+Moreover, to fully realize acceleration in long-context processing, KV cache compression algorithms must be carefully designed.
+Previous works [13, 14, 15] have mainly focused on compressing the KV cache for the generation stage only, and this approach fails to reduce prefill latency (Figure 1(b)).
+Therefore, to enable efficient long-context processing, it is crucial to develop methods that not only compress the KV cache but also accelerate the computation in the prefill stage.
+2.2 KV Cache Compression
 Figure 2: Comparison of the number of tokens processed in each layer/head of LLMs during prefill computation and KV caching across different KV cache compression techniques.
@@
 The blue background box indicates the set of sharing selected token indices.
-Recent advancements in large language models (LLMs) have enabled the processing of long-context sequences, such as those comprising 128k tokens Achiam et al.
-(2023); Gemini Team et al.
-(2023); Anthropic .
-This capability significantly broadens the range of applications for LLMs Yi et al.
-(2024); Laban et al.
-(2023); Gu (2023).
-However, as the length of the input sequence increases, the size of these key-value (KV) caches also increases, making them a significant bottleneck in the serving of long-context LLMs.
-Therefore, compression of KV caches is essential for optimizing the operation of LLMs.
-There has been active research on compressing KV cache to alleviate its burden in long-context handling Zhang et al.
-(2023); Xiao et al.
-(2023; 2024); Oren et al.
-(2024); Chen et al.
-(2024).
-Some techniques have demonstrated the ability to preserve accuracy comparable to full-context processing after the compression Li et al.
-(2024); Feng et al.
-(2024); Fu et al.
-(2024), while there exists a compression technique that has achieved inference speedup in both the prefill and generation stages Shi et al.
-(2024).
-However, no KV cache compression techniques currently exist that can simultaneously preserve accuracy and achieve inference speedup, particularly during the prefill stage.
-In this paper, we propose FastKV, an innovative approach designed to expedite long-context handling with LLMs while maintaining their accuracy.
-FastKV stems from our finding that the properties of attention maps differ between the early and later layers of LLMs.
-Through detailed analysis of LLMs, we discover that in the later layers, attention maps concentrate on a limited set of significant tokens, which remain consistent across layers.
-In contrast, the early layers engage with a broader array of tokens and exhibit diverse attention patterns as they process the context of the input prmopt.
-Based on these findings, FastKV adopts a novel Token-Selective Propagation (TSP) method that applies different strategies of KV cache compression at the early and later layers respectively As shown in Figure 1, this dual-strategy approach enables FastKV to improve both time-to-first-token (TTFT) and throughput of long-context processing while effectively preserves the accuracy of LLMs.
-Our experimental results show that FastKV can achieve 2.00$\times$ speedup for TTFT and 1.40$\times$ improvement for throughput compared to HeadKV, while maintaining a similar level of accuracy with less than a 1% accuracy gap.
-These results demonstrate that FastKV promises a practical solution for long context scenarios, particularly in real-time applications that require efficient KV cache management and low latency for the prefill and generation stages.
-2 Background
-2.1 Long-Context Processing with LLMs
-LLMs have significantly enhanced the natural language processing (NLP) abilities of AI systems Brown et al.
-(2020); Radford (2018); Kamalloo et al.
-(2023); Jiang et al.
-(2021) through the use of attention mechanisms Vaswani (2017).
-These attention mechanisms carefully evaluate the relationships between tokens within sentences to construct context vectors by aggregating token data according to the attention scores to extract context-specific information Brauwers & Frasincar (2021).
-Distinct from earlier NLP models such as RNNs Sherstinsky (2020) and LSTMs Hochreiter (1997), attention mechanisms consider the entire token data from input sequences, thereby avoiding issues of prompt forgetting.
-Benefiting from these advancements, recent researches indicates that LLMs are capable of processing long-context sequences.
-However, the very nature of LLMs, which retain all token data to facilitate attention mechanisms, leads to significant memory overhead.
-Typically, token data for attention mechanisms are stored in the form of KV cache, and the size of KV cache escalates with sequence length, becoming a primary source of inefficiency in LLMs when processing long-context scenarios.
-For instance, in the LLaMA-3.1-8B model Dubey et al.
-(2024), managing a sequence of 128k tokens requires 17.12GB of memory for KV cache, whereas the model itself only requires 16.06GB.
-Therefore, the excessive KV cache demands form a critical bottleneck in LLM inference during long-context processing, and efficient compression of KV cache is crucial for streamlined LLM operation.
+KV cache compression techniques for long-context processing are primarily designed to prune tokens, thereby reducing the associated KV caches and effectively managing memory usage.
+A common approach leverages the attention scores generated by each attention head to selectively prune tokens in an attention-head-wise manner.
+This method operates on the premise that token importance varies across attention heads.
+A notable example is SnapKV in Figure 2(a).
+It discovers that each attention head in the model consistently focuses on specific prompt token during generation, and this robust pattern can be obtained from an observation window located at the end of the input prompt.
+Thus, SnapKV evaluates the token importance score for each attention head by summing the attention scores assigned to each token as follows:
+$S^{l,h}_{i}=\frac{1}{2w_{p}+1}\sum_{m=-w_{p}}^{w_{p}}\sum_{n=0}^{N_{\text{obs}}}Att_{l}[h,\,N_{I}-n,\,i+m]$ (1)
+Here, $S_{i}^{l,h}$ is the importance score of $i$-th token in $h$-th attention head of the $l$-th layer.
+$Att_{l}$ denotes the attention score matrix of $l$-th layer, while $N_{I}$ and $N_{obs}$ indicate the number of tokens in the input prompt and the observation window, respectively.
+SnapKV applies average pooling on the scores to prevent sparse token selection.
+Here, $12w_{p}+$ 1 indicates the size of pooling window.
+SnapKV selects tokens from the observation window first and then selects tokens with the highest importance scores for KV caching.
+Since SnapKV relies on the attention scores produced by each attention head, it requires propagating the full input prompt throughout the entire model, during the prefill stage.
+As a result, SnapKV cannot improve computational efficiency or reduce the latency of the prefill stage, which directly impacts TTFT, a key factor in optimizing user experience for LLM serving.
+AdaKV and HeadKV in Figure 2(b)(c) extend SnapKV by applying dynamic KV budget allocation to head groups [22] and individual heads, respectively, which improves accuracy under extreme compression such as when the average KV budget is 128.
+However, when applied to more realistic KV budget targets, such as 512, 1024 and 2048, dynamic KV budget allocation exhibits marginal accuracy improvements or even a slight accuracy degradation compared to SnapKV, while they introduce notable latency and throughput overhead.
+The primary reason of the overhead is that the processing speed of the attention mechanism is limited by the attention head with the longest KV cache.
+As the size of KV budget increases, the imbalance in KV cache sizes across attention heads becomes more pronounced and it amplifies latency overhead.
+This overhead makes these methods less practical for real-world applications, where achieving low latency is critical.
+In contrast, GemFilter in Figure 2(d) adopts a fundamentally different approach that enhances the computational efficiency of the prefill stage.
+This method identifies relevant tokens at the layer level in intermediate layers, called the filter layer, and uses these indices to compress the input prompt.
+Then, the prefill stage is recomputed with compressed input prompt.
+By processing full-context with part of the LLM, GemFilter effectively reduces TTFT.
+However, as the removal of tokens from the input prompt results in the complete loss of any information embedded in the discarded tokens, GemFilter struggles to maintain general performance on complex tasks or tasks requiring comprehensive contextual understanding, even when given moderate KV budgets such as 2K.
+Thus, there are currently no KV cache compression techniques that can adequately preserve accuracy while simultaneously improving TTFT and throughput.
+This underscores the need for new methods that effectively balance these objectives.
+3 Proposed FastKV
+3.1 Overview of FastKV
+Figure 3: Illustration of the proposed FastKV scheme.
+The proposed FastKV introduce Token-Selective Propagation approach to selectively propagate only a limited set of tokens while effectively compressing KV data.
+GemFilter accelerates the prefill stage while performing effective KV cache compression.
+However, by removing a portion of the tokens from the input prompt, it inevitably causes complete loss of the discarded information.
+Critically, LLMs rely on attention mechanisms to enable rich interactions across tokens, and GemFilter fails to exploit this characteristic.
+For instance, GemFilter performs well on multi-document QA tasks in the LongBench evaluation, which primarily require information retrieval, but exhibits substantially lower performance on tasks requiring comprehensive context understanding, such as code completion.
+To overcome this limitation, we propose FastKV, a KV cache compression framework that accelerates the prefill stage while preserving global context.
+The overview of FastKV is illustrated in Figure 3.
+To speed up the prefill stage, FastKV employs the Token-Selective Propagation (TSP) designed to maintain full-context processing in early layers and selectively propagates only a small set of semantically important tokens to later layers.
+The TSP method identifies a subset of crucial tokens (e.g., 2K) at a mid-layer (TSP layer) and propagates only these tokens through subsequent layers.
+These selected tokens, having already aggregated contextual information through the attention mechanisms in the early layers, encapsulate the essential semantics of the input sequence and thus preserve global context throughout the later layers.
+Furthermore, the TSP approach reduces the number of tokens processed in the later layers, thereby accelerating the processing speed in these layers.
+To perform KV cache compression, FastKV evaluates the attention scores generated by each attention head group to determine the importance of KV data.
+Based on the head-wise scores $S_{i}^{j,h}$ (Equation 1), the group-level importance is computed as follows:
+$S_{i}^{l,g}=\frac{1}{H_{G}}\sum_{h=h_{g}}^{h_{g}+H_{G}}S_{i}^{j,h}$ (2)
+Here, $H_{G}$ denotes the number of attention heads in each group.
+$g$ and $h_{g}$ denotes the group index and the index of the first head in the group.
+While FastKV applies the same KV cache compression technique to both early and later layers, in the later layers, compression is performed over a reduced set of KV data corresponding to the tokens selected by the TSP approach, allowing for more precise identification of important information.
+It is important to note that KV data compression occurs solely during the storing of KV data to establish KV caches for input prompt.
+Consequently, FastKV ensures that the entire set of KV data, including those of the removed tokens, is propagated to the subsequent layer until the TSP layer.
+This approach guarantees full-context propagation to preserve essential contextual information from the input prompt.
+This dual strategy enables FastKV to achieve significant speed improvements without compromising accuracy, as it properly preserves the contextual information of input sequences.
+3.2 Token-Selective Propagation
+TSP improves prefill efficiency by propagating only a subset of important tokens while retaining critical contextual information.
+It is defined by three factors: the layer for token selection (TSP layer), identification of important tokens and the number of tokens propagated (TSP length).
+At the selected TSP layer, token importance is assessed by averaging attention scores across all heads as follows:
+$S_{i}^{TSPlayer}=\frac{1}{H}\sum_{h=0}^{H-1}S_{i}^{TSPlayer,h}$ (3)
+Here, $H$ denotes the number of attention heads in each layer.
+Since compressing the layer output requires a single index set, we average the attention weights to calculate a score $S_{i}^{TSPlayer}$ that represents the importance of tokens at the layer level, as outlined in Equation 3.
+The top-k tokens based on $S_{i}^{TSPlayer}$ are selected, and their indices are combined with window tokens to form the final set of TSP tokens indices propagated to the next layer.
+These tokens are used to index hidden states after the MLP block, effectively reducing the layer input size for subsequent computation.
+Unlike GemFilter, which restarts computation from the first layer using only selected tokens, TSP continues inference from the TSP layer without recomputation.
+Moreover, whereas prior methods retain hidden states for all tokens through the prefill stage, TSP discards low-importance tokens after the TSP layer, reducing computation while preserving contextual information.
+To validate this design, we further analyze the effect of decoupling TSP from the KV budget in Section 3.3 and how TSP maintains global semantics despite this token reduction in Section 3.4.
+3.3 TSP Length and KV Budget
+Figure 4: (a) Minimum match rate between the 2K important tokens selected at each layer from 128K tokens and those selected by its subsequent layers.
+(b) Comparison of normalized L2 distances between hidden states generated by the full-context baseline, FastKV, and GemFilter-like methods.
+The GemFilter-like method represents the approach where the input prompt is pruned by retaining only selected tokens and inference is recomputed, following the strategy of GemFilter.
+As shown in Figure 4(a), we measure the minimum match rate as the lowest overlap ratio between the top-2K important tokens selected at a given layer from 128K input tokens and those selected by all subsequent layers.
+The match rate of the important tokens is around 25%.
+This rate is sufficiently high to justify the use of the TSP approach with an adequate TSP length (e.g.
+2K).
+However, a short TSP length (e.g.
+512) may lead to significant information loss.
+Therefore, TSP length, which defines the number of tokens selected for propagation, is strategically designed to operate independently of the KV cache budget.
+For instance, if the KV cache budget is set to 512 while the TSP length is 2048, TSP propagates 2048 tokens during its operation, but only the final 512 tokens are stored in the KV cache.
+This separation allows TSP to process a significantly larger set of tokens during propagation, enabling it to capture more contextual information while maintaining the KV cache within its predefined budget.
+As shown in Section 4.4, a TSP length of 2048 has been found sufficient to preserve accuracy of the model under specific KV budget while maintaining efficiency in terms of TTFT.
+3.4 Impact of TSP on LLM Output
+While FastKV is designed to accelerate the prefill stage while preserving the model’s ability to process inputs in full-context, it is critical to empirically verify whether TSP effectively retains global contextual information.
+In this section, we quantitatively evaluate the impact of TSP on the distribution of LLM hidden states compared to a GemFilter-like approach.
+As shown in Figure 4(b), we use a sampled needle-in-a-haystack data for 128K input prompt and identify the top-2K important tokens at a target layer based on attention scores computed via Equation 3.
+We perform inference over the 128K input prompt for the full-context baseline, indexing the output hidden states corresponding to the selected tokens.
+For the GemFilter-like variant, we prune the input prompt by keeping only the selected tokens and recompute inference.
+For the FastKV, we apply TSP at the target layer, propagate only the selected tokens through subsequent layers.
+We then measure the L2 distance between the hidden states obtained from the full-context inference and those obtained from FastKV and GemFilter, respectively, using the Mean Squared Error (MSE) metric.
+Figure 4(b) shows the normalized L2 distances across layers.
+For 2 different models, FastKV consistently achieves substantially lower L2 distances compared to the GemFilter-like variant, particularly from the middle layers onward.
+This indicates that TSP effectively preserves the global structure of hidden states, maintaining a close approximation to the full-context model despite aggressive token reduction.
+In contrast, the GemFilter-like approach exhibits consistently large L2 distances across layers, suggesting that prompt-level token pruning causes an irreversible loss of contextual information.
+These findings highlight the advantage of applying token selection after early-layer attention mixing, as in TSP, which enables the model to retain essential semantic information across the sequence.
+Thus, FastKV successfully balances prefill acceleration and contextual fidelity, achieving compression without sacrificing the model’s ability to capture global context.
+3.5 TSP Layer Selection
+Building on the analysis in Section 3.4, we describe the method for selecting an appropriate TSP layer to balance the trade-off between acceleration and performance in LLM inference.
+Unlike GemFilter, which selects a layer based on its ability to attend to specific tokens (e.g., a "needle" token), FastKV selects the TSP layer based on how effectively the compressed computation at that layer preserves the output distribution of the original full-context model.
+Specifically, we compute the MSE between the hidden states of the full-context model $\mathbf{H}$ and those of the TSP-compressed model $\mathbf{H^{\prime}_{L}}$ at various candidate layers $L$, and choose the layer that minimizes this discrepancy, as defined by the following objective:
+$L_{TSP}=\operatorname*{argmin}_{L\leq L_{\max}}\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{H}_{i}-\mathbf{H}^{\prime}_{L,i}\right|_{2}^{2}$ (4)
+Here, $N$ is the number of tokens used for TSP indices and $L_{max}$ denotes the maximum allowable TSP depth, set to half the total number of layers, to ensure sufficient opportunity for acceleration without applying TSP to overly later layers.
+This strategy ensures that TSP is applied at a point where the model has already integrated sufficient global context, enabling downstream layers to benefit from compression without degrading semantic fidelity.
+4 Experiments
+Table 1: LongBench results comparison between baseline methods and FastKV.
+Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Synthetic
+Method NrtvQA Qasper MF-en HotpotQA 2WikiMQA MuSiQue GovReport QMSum MultiNews TREC TriviaQA SAMSum LCC RB-P PCount PRe Avg.
+LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = Full
+Full KV 23.59 39.33 51.80 56.89 45.53 34.66 33.57 22.96 26.10 73.50 90.88 42.49 62.03 67.05 2.00 89.50 47.62
+LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = 512
+SnapKV 24.53 36.47 50.26 55.91 45.49 34.98 26.67 22.69 24.34 73.00 90.05 41.57 60.61 64.01 2.00 89.00 46.35
+AdaKV 23.79 36.56 50.77 56.37 45.08 34.48 26.97 22.93 24.19 73.00 90.80 42.01 61.32 65.83 2.00 88.50 46.54
+HeadKV 23.76 37.88 51.19 55.65 45.64 34.93 29.73 22.94 26.01 73.00 90.71 40.28 62.54 66.19 2.00 88.50 46.93
+GemFilter 16.44 26.31 46.58 47.38 42.32 27.87 26.21 19.42 22.33 70.00 78.17 37.20 29.98 38.39 2.00 44.00 35.91
+FastKV 23.34 36.35 51.84 56.29 46.44 34.11 26.13 22.83 24.10 74.00 90.50 41.24 60.98 65.22 2.00 87.50 46.43
+LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = 2048
+SnapKV 22.86 38.14 52.41 56.49 45.98 34.82 31.46 22.89 25.97 73.50 91.00 42.18 62.52 67.12 2.00 89.50 47.43
+AdaKV 22.97 38.67 52.36 56.46 44.98 34.63 31.04 23.22 25.80 73.50 90.99 42.45 62.12 67.49 2.00 89.50 47.39
+HeadKV 22.47 38.84 52.18 55.62 45.14 34.84 32.20 23.10 26.08 73.50 90.99 41.80 62.65 67.02 2.00 89.00 47.34
+GemFilter 21.22 35.02 53.58 56.37 47.02 31.29 31.36 20.72 25.62 73.50 89.86 39.82 48.54 52.56 1.50 92.00 45.00
+FastKV 23.65 38.49 52.98 56.80 46.83 34.23 30.83 23.05 25.76 75.50 90.99 42.72 62.00 67.14 2.00 87.50 47.53
+Mistral-Nemo-12B-Instruct, KV Budget = Full
+Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 1.50 98.00 48.51
+Mistral-Nemo-12B-Instruct, KV Budget = 512
+SnapKV 23.58 40.12 55.17 47.91 45.69 25.17 23.77 22.48 24.08 74.00 89.44 43.09 68.31 62.15 1.00 97.50 46.47
+AdaKV 24.19 39.65 54.30 48.16 45.43 26.60 23.06 22.85 23.31 74.00 89.26 43.09 67.38 62.41 1.50 98.50 46.48
+HeadKV 25.88 40.57 57.12 48.11 46.11 25.26 26.19 22.92 25.25 74.50 89.74 41.81 68.79 64.76 1.00 97.50 47.22
+GemFilter 25.45 37.64 53.91 52.83 51.26 33.02 26.23 19.42 23.87 65.50 84.16 40.05 38.03 41.13 4.00 57.50 40.88
+FastKV 25.51 41.48 56.76 48.73 45.43 26.54 23.58 22.20 24.00 74.00 89.42 43.22 70.23 65.55 1.00 98.00 47.23
+Mistral-Nemo-12B-Instruct, KV Budget = 2048
+SnapKV 24.35 42.95 56.92 49.08 46.43 26.26 28.27 23.68 26.07 75.00 89.82 44.20 68.89 68.00 1.00 97.50 48.03
+AdaKV 24.67 42.46 56.05 48.60 45.97 26.23 26.58 23.39 25.74 75.00 89.66 44.09 68.25 67.70 1.50 97.00 47.68
+HeadKV 26.05 42.15 57.46 48.49 45.85 25.86 30.27 24.22 26.17 75.00 89.66 43.74 68.64 67.29 1.50 98.50 48.18
+GemFilter 26.42 42.40 56.98 57.64 53.92 34.33 30.51 21.60 25.96 72.00 89.65 44.48 48.34 48.06 3.50 97.00 47.05
+FastKV 26.61 43.37 57.12 49.41 47.22 26.27 28.40 23.27 25.81 75.00 89.72 45.04 70.28 69.24 0.50 98.00 48.53
+4.1 Setup
+Models and Datasets.
+We evaluate two open-source LLMs of different sizes: LLaMA-3.1-8B-UltraLong-1M-Instruct [8] and Mistral-Nemo-12B-Instruct [20].
+Built on LLaMA-3.1-8B-Instruct [7], LLaMA-3.1-8B-UltraLong-1M-Instruct is a model trained to support 1M context window size with 32 decoder layers.
+Mistral-Nemo-12B-Instruct has 40 decoder layers and employs a context window size of 128K tokens.
+To assess long-context understanding and retrieval capabilities of the LLMs, we use two benchmark datasets: LongBench [17] and Needle-in-a-Haystack [18].
+See Appendix B for more details.
+Implementation Details.
+We integrate our proposed FastKV method upon self-attention implementation of HuggingFace Transformers library, which utilizes FlashAttention-2 [19] kernel.
+We select layer 15 as the TSP layer for LLaMA-3.1-8B-UltraLong-1M-Instruct and layer 19 for Mistral-Nemo-12B-Instruct.
+We fix the observation window size ($N_{obs}$) to 8 and the kernel size of average pooling ($12w_{p}+$ 1) to 7.
+The TSP lengths are set to 2048 for evaluation.
+Baselines.
+We compare our approach with four baseline methods for KV cache compression: SnapKV [13], AdaKV [14], HeadKV [15], and GemFilter [16].
+To ensure a fair comparison, SnapKV, AdaKV, and HeadKV use the same local window size, pooling method, and pooling kernel size as FastKV.
+The indices of GemFilter filter layers for LLaMA-3.1-8B-UltraLong-1M-Instruct and Mistral-Nemo-12B-Instruct are 13 and 19, respectively [16].
+4.2 Accuracy Evaluation
+LongBench.
+The accuracy evaluation results on LongBench are summarized in Table 1.
+On LLaMA-3.1-8B-UltraLong-1M-Instruct, previous works such as SnapKV, AdaKV, and HeadKV, which employ token removal after prefilling all the tokens at each layer, successfully maintain accuracy after KV cache compression, with an average accuracy drop of less than 0.69% and 0.19% compared to Full KV for 512 and 2048 KV budgets, respectively.
+GemFilter, which recomputes the prefill stage with the fragmented input prompt which consists only of selected tokens, completely loses the information of the discarded tokens.
+This leads to a significantly larger accuracy drop, with reductions of up to 11.71% and 2.62% compared to Full KV for 512 and 2048 KV budgets, respectively.
+On the other hand, the proposed FastKV preserves accuracy on par with Full KV and other methods such as AdaKV and HeadKV.
+Moreover, on Mistral-Nemo-12B-Instruct, FastKV outperforms other methods and consistently achieves strong accuracy across LongBench, while achieving memory and computational efficiency comparable to GemFilter.
+The advantage of FastKV is especially pronounced in code completion tasks, which demand a comprehensive understanding of global context.
+Figure 5: Needle-in-a-Haystack results of LLaMA-3.1-8B-UltraLong-1M-Instruct with 2048 KV budget.
+Figure 6: Latency and throughput evaluation results of LLaMA-3.1-8B-UltraLong-1M-Instruct on a single NVIDIA A100 GPU.
+KV budgets and TSP length are both set to 2048 throughout the experiments.
+(a) TTFT with context length of 128K and 256K tokens.
+(b) Token throughput with context length 128K.
+Needle-in-a-Haystack.
+We present the Needle-in-a-Haystack (NIAH) evaluation results for LLaMA-3.1-8B-UltraLong-1M-Instruct with a KV budget of 2048 in Figure 5.
+The evaluation is conducted with input context lengths up to 256K tokens.
+Additional results for other LLMs and KV budgets are provided in Appendix C.2.
+Notably, FastKV achieves performance comparable to other methods even under 256K long-context inputs.
+Taken together with the LongBench evaluation, this results demonstrate that FastKV effectively maintains general-purpose LLM performance across diverse long-context tasks while achieving efficient KV cache compression and inference acceleration.
+4.3 Efficiency Evaluation
+We evaluate the time-to-first-token (TTFT) of the prefill stage, throughput of token generation and memory usage of KV cache with LLaMA-3.1-8B-UltraLong-1M-Instruct on a single NVIDIA A100 GPU.
+Figure 6 reports TTFT results with 128K/256K input, and throughput results with 128K input and 2048 KV budget.
+More detailed evaluation results can be found in Appendix C.3.
+TTFT.
+As shown in Figure 6(a), Both GemFilter and FastKV achieve significant TTFT improvements over other baselines by reducing the computational complexity of the prefill stage through the concept of propagating only selected tokens during this stage.
+As a result, FastKV achieves TTFT that is 1.97$\times$ shorter than Full KV with 128K input and 1.99$\times$ shorter than HeadKV with 256K input.
+While both methods enhance efficiency, GemFilter incurs additional overhead due to the need for prefill recomputation, whereas FastKV introduces extra processes for compressing generated KV caches.
+These two factors offset each other.
+Consequently, GemFilter’s slight advantage in TTFT is primarily attributed to its earlier filter layer (13) compared to the TSP layer (15) for LLaMA-3.1-8B-UltraLong-1M-Instruct.
+In contrast, SnapKV, AdaKV, and HeadKV exhibit even longer TTFT than Full KV.
+This is because they process full-context information during the prefill stage, similar to Full KV, while also incurring additional overhead from indexing and selecting crucial tokens.
+Throughput.
+We measure the throughput for generating a total of 128 tokens.
+As shown in Figure 6(b), all KV compression techniques result in throughput improvements compared to Full KV.
+Both FastKV and GemFilter achieve a significant level of throughput improvement.
+FastKV’s throughput reaches 4.82$\times$ of Full KV and 1.97$\times$ of HeadKV with 2048 KV budget.
+In contrast, SnapKV, which stores KV caches of different tokens set for each attention head, shows slightly lower throughput.
+Additionally, HeadKV, which allocates varying KV budgets across attention heads, exhibits the lowest throughput among KV cache compression techniques.
+Managing different sequence lengths for keys and values across attention heads increases the complexity of the attention mechanism.
+AdaKV also adopts fine-grained KV budget allocation similar to HeadKV except that its granularity is attention-group.
+AdaKV mitigates the influence of fragmented KV budget to some extent thanks to GQA-aware compression, but it does not achieve the same level of throughput with FastKV and GemFilter.
+Figure 7: Comparison across different KV cache compression methods
+Memory Usage.
+As shown in Table 2, FastKV, GemFilter, and AdaKV reduce KV cache memory usage by 31.81$\times$ over the Full KV baseline for 256K tokens by retaining only a selected subset of tokens.
+This is achieved by maintaining shared token indices per KV head group, unlike SnapKV and HeadKV which store separate indices for each attention head.
+A comprehensive comparison of latency, throughput, and accuracy across all methods is summarized in Figure 7, highlighting FastKV as the only approach that achieves favorable performance across all metrics.
+Table 2: KV cache memory usage for a 256K tokens in LLaMA-3.1-8B-UltraLong-1M-Instruct. KV budget is set to 2048 for each compression method.
+Method Full KV SnapKV AdaKV HeadKV GemFilter FastKV
+KV Cache Size (GB) 34.36 4.30 1.08 4.30 1.08 1.08
+Figure 8: LongBench results for LLaMA-3.1-8B-UltraLong-1M-Instruct. (a) Effect of TSP layer index on accuracy and TTFT (KV budget = 2048). (b) Effect of TSP length across KV budgets.
+4.4 Ablation Studies
+TSP Layer Index.
+We evaluate the impact of the TSP layer index on model accuracy under a fixed KV budget of 2048.
+As shown in Figure 8(a), applying TSP at earlier layers degrades accuracy due to insufficient context mixing, while performance improves as the TSP layer moves deeper.
+The proposed selection algorithm identifies layer 15 as an optimal point where the LongBench score begins to saturate, indicating a favorable trade-off between speed and accuracy.
+TSP Length.
+We investigate the impact of TSP length on model performance under varying KV budgets.
+TSP length and KV budget control acceleration in the prefill and generation stages, respectively, and are independently tunable.
+As shown in Figure 8(b), accuracy drops noticeably when the TSP length is below 2048, while it saturates beyond that point.
+Since shorter lengths reduce latency but limit context propagation, a TSP length of 2048 provides an effective balance between latency and accuracy.
+This configuration is adopted in our main experiments to optimize prefill efficiency without compromising general performance.
+5 Conclusion
+In this paper, we present FastKV, a novel KV cache compression framework that enhances the efficiency of long-context LLM inference while preserving model accuracy.
+FastKV introduces Token-Selective Propagation (TSP), a mechanism that allows the early layers of the model to attend to the full input sequence, while the later layers operate only on a reduced set of semantically important tokens.
+This design significantly reduces the computational burden of the prefill stage without discarding critical contextual information.
+Additionally, FastKV compresses the KV cache for the selected tokens, enabling additional efficiency in memory usage without introducing latency bottlenecks.
+Experimental results show that FastKV achieves higher throughput and lower TTFT than baseline methods while maintaining high accuracy on long-context benchmarks.
+Appendix A Additional Methodological Details
+A.1 FastKV Algorithm
+We present the pseudocode for the FastKV algorithm, which performs two complementary compression strategies during the prefill stage: (1) reducing hidden states via Token-Selective Propagation (TSP), and (2) compressing the attention key-value (KV) cache at each layer.
+These two steps are carried out independently: TSP selects a subset of hidden states to propagate to later layers, while KV cache compression prunes KV cache entries at each layer using attention-based importance scores.
+Together, they improve prefill efficiency and reduce memory usage without modifying the model architecture or introducing any runtime overhead during decoding.
+The key terms used in the pseudocode are:
+  • LTSP: The index of TSP layer
+  • BTSP: TSP length
+  • BKV: KV budget
+  • KVCompress: Selects the KV entries of top-BKV critical tokens as determined by Equation 2.
+  • HiddenCompress: Selects hidden states of top-BTSP critical tokens, as determined by Equation 3, to propagate to next layer.
 Algorithm 1 FastKV algorithm for the KV cache compression during the prefill stage
-1:$input$ $sequence$$\{I\},$$\#layers$$\{L\},$ $TSPlayer,$ $TSPlength$$\{B_{TSP}\},$ $KVbudget$$\{B_{KV}\}$
+1:$input$ $sequence$$\{I\},$$\#layers$$\{L\},$ $L_{TSP},$ $B_{TSP},$ $B_{KV}$
 2:$generated$ $token$$\{O\},$ $KV$ $Cache$$\{C\}$
@@
 4:for $l$ = 0 to $L-1$ do
-5:     if $l\leq TSPlayer$ then
-6:         $X,Att_{l},K_{X},V_{X}\leftarrow layer_{l}(X)$
+5:     if $l\leq L_{TSP}$ then
+6:         $X,Att_{l},K_{X},V_{X}\leftarrow layer_{l}(X)$triangleright $X:$ $hidden$ $states$ $before$ $TSP$
 7:         $K,V\leftarrow KVCompress(K_{X},V_{X},Att_{l},B_{KV})$
-8:         if $l==TSPlayer$ then
-9:              $x\leftarrow HiddenCompress(X,Att_{l},B_{TSP})$
+8:         if $l==L_{TSP}$ then
+9:              $x\leftarrow HiddenCompress(X,Att_{l},B_{TSP})$triangleright $x:$ $reduced$ $hidden$ $states$
 10:         end if
@@
 18:return $O,C$
-2.2 KV Cache Compression
-KV cache compression techniques for long-context processing are primarily designed to prune tokens, thereby reducing the associated KV caches and effectively managing memory usage.
-A common approach leverages the attention scores generated by each attention head to selectively prune tokens in an attention-head-wise manner.
-This method operates on the premise that token importance varies across attention heads.
-A notable example is SnapKV Li et al.
-(2024) in Figure 2(a).
-It discovers that each attention head in the model consistently focuses on specific prompt token during generation, and this robust pattern can be obtained from an observation window located at the end of the input prompt.
-Thus, SnapKV evaluates token importance score for each attention head by summing the attention scores assigned to each token as follows:
-$S^{l,h}_{i}=\frac{1}{2w_{p}+1}\sum_{m=-w_{p}}^{w_{p}}\sum_{n=0}^{N_{\text{obs}}}Att_{l}[h,\,N_{I}-n,\,i+m]$ (1)
-Here, $S_{i}^{l,h}$ is the importance score of $i$-th token in $h$-th attention head of the $l$-th layer.
-$Att_{l}$ denotes the attention score matrix of $l$-th layer, while $N_{I}$ and $N_{obs}$ indicate the number of tokens in the input prompt and the observation window, respectively.
-SnapKV applies average pooling on the scores to prevent sparse token selection.
-Here, $12w_{p}+$ 1 indicates the size of pooling window.
-SnapKV selects tokens from the observation window first and then selects tokens with the highest importance scores for KV caching.
-Since SnapKV relies on the attention scores produced by each attention head, it requires propagationg of the full input prompt throughout the entire model, during the prefill stage.
-As a result, SnapKV cannot improve computational efficiency or reduce the latency of the prefill stage, which directly impacts TTFT, a key factor in optimizing user experience for LLM serving.
-AdaKV Feng et al.
-(2024) in Figure 2(b) expands SnapKV approach to grouped-query-attention (GQA) mechanism Ainslie et al.
-(2023) by introducing dynamic KV budget allocation across attention groups and selecting important tokens in attention-group-wise manner.
-To enhance accuracy preservation, HeadKV Fu et al.
-(2024) in Figure 2(c) builds upon SnapKV by introducing dynamic KV budget allocation across attention heads.
-This feature has shown effective in preserving accuracy under aggressive KV cache compression scenarios, such as when the average KV budget is as low as 128.
-However, when applied to more realistic KV budget targets, such as 512, dynamic KV budget allocation exhibits marginal accuracy improvements or even a slight accuracy degradation compared to SnapKV, while they introduces notable latency and throughput overhead.
-The primary reason of the overhead is that the processing speed of the attention mechanism is limited by the attention head with the longest KV cache.
-As the size of KV budget increases, the imbalance in KV cache sizes across attention heads becomes more pronounced and it amplifies latency overhead.
-This overhead makes these methods less practical for real-world applications, where achieving low latency is critical.
-In contrast, GemFilter Shi et al.
-(2024) in Figure 2(d) adopts a fundamentally different approach that enhances the computational efficiency of the prefill stage.
-GemFilter is based on the observation that LLMs can identify relevant tokens in intermediate layers, referred to as the filter layer.
-The filter layer selects the indices of relevant tokens at the layer level rather then head level, and GemFilter uses these indices to compress the input prompt.
-Then, the prefill stage is recomputed with compressed input prompt.
-By processing full-context with part of the LLM, GemFilter effectively reduces TTFT.
-The GemFilter paper reports successful accuracy preservation when using a KV cache budget of 2k, where the input prompt is compressed with 2k tokens.
-Moreover, as GemFilter focuses on the selected tokens more intensively than a full-context model by only processing selected tokens, with sufficient KV budget, there are some cases where it achieves even better accuracy than the full-context model.
-However, as the removal of tokens from the input prompt results in the complete loss of any information embedded in the discarded tokens, GemFilter struggles to maintain accuracy with lower KV budgets, such as 512, and even with 2k KV budgets when handling complex tasks.
-Thus, there are currently no KV cache compression techniques that can adequately preserve accuracy while simultaneously improving TTFT/throughput.
-This underscores the need for new methods that effectively balance these objectives.
-3 Proposed FastKV
-3.1 Overview of FastKV
-The overview of FastKV is illustrated in Figure 2(e) and Algorithm 1.
-FastKV implements distinct KV cache compression strategies for the early and later layers of LLMs.
-To speedup the prefill stage, FastKV employs the TSP approach in the later layers.
-This approach is based on the insight that after the early layers have analyzed contextual information, the later layers tend to focus consistently on the same crucial tokens.
-The TSP approach, therefore, involves selectively propagating only a limited set of tokens (e.g.
-2k tokens) identified by the TSP layer, which is strategically positioned in the middle of the LLM architecture.
-The TSP layer utilizes its attention map to pinpoint these critical tokens.
-Unlike the early layers, which propagate the entire set of tokens, the TSP approach in the later layers transmits only the selected tokens to subsequent layers.
-This targeted propagation means that the later layers process far fewer tokens, thus it can accelerate the processing speed in these layers.
-While the TSP approach results in a smaller set of KV data for the selected tokens, FastKV still applies the same KV cache compression techniques used in the early layers to optimize the compression ratio.
-In the early layers, FastKV evaluates the attention scores generated by each attention head to determine the importance of KV data.
-GQA compatible KV cache compression of FastKV selectively removes the KV data of tokens that receive low attention scores in an attention-group-wise manner.
-It is important to note that KV data compression occurs solely during the saving of KV data to establish KV caches for input prompt.
-Consequently, FastKV ensures that the entire set of KV data, including those of the removed tokens, is propagated to the subsequent layer until the TSP layer.
-This approach guarantees full context propagation to preserve essential contextual information from the input prompt.
-This dual strategy enables FastKV to achieve significant speed improvements without compromising accuracy, as it properly preserves the contextual information of input sequences.
-3.2 Proposed TSP
-TSP aims to enhance the efficiency of the prefill stage by propagating only a selected set of important tokens, similar to the approach used by GemFilter.
-However, in response to the insights obtained from the previous sections, TSP is designed to retain the necessary scope of token propagation to ensure that critical contextual information is preserved.
-The operation of TSP is governed by three key factors: the layer for token selection (TSP layer), identification of important tokens, and the number of tokens selected for propagation (TSP length).
-TSP begins by identifying TSP length important tokens from the full tokens at the TSP layer.
-The method for identifying important tokens in TSP builds upon Equation 1.
-The propagation of selected tokens across LLM layers involves transmitting the hidden states associated with those selected tokens.
-Consequently, TSP necessitates the assessment of token importance on a layer-by-layer basis.
-To facilitate this, TSP includes an additional step that aggregates the attention scores from all attention heads within the TSP layer as follows:
-$S_{i}^{TSPlayer}=\frac{1}{H}\sum_{h=0}^{H-1}S_{i}^{TSPlayer,h}$ (2)
-Here, $H$ denotes the number of attention head in each layer.
-From the TSP layer onward, TSP propagates only selected important tokens through the later layers.
-Unlike GemFilter, which reverts to the first layer of the LLM with selected tokens and recomputes the prefill stage, TSP directly propagates the selected tokens to subsequent layers without returning to the first layer.
-Meanwhile, although previous approaches to KV cache compression aimed at high accuracy by propagating the entire context throughout the prefill stage, the proposed TSP method discards the unimportant tokens after the TSP layer.
-Therefore, the implementation of TSP must be carefully designed based on a detailed analysis of the properties of token importance and its impact on the LLM output.
-3.3 Impact of Important Tokens
-Figure 3: Rate of sum of attention scores for the top 2k important tokens selected to the total attention score at each layer.
-In the LLM, the attention mechanism is the critical component for extracting contextual information, while other parts of the model process input in a more token-isolated manner.
-Within the attention mechanism, each token generates a new embedding by integrating the embeddings of all preceding tokens, with this integration weighted by the attention scores.
-Consequently, tokens that receive low attention scores have only a marginal impact on the processing of contextual information.
-To ensure the removal of the unimportant tokens without significant impact on the LLM output, the important tokens must receive significantly higher attention scores.
-We evaluate the actual impact of these important tokens by examining their attention scores relative to those of the remaining tokens, as shown in Figure 3.
-In this analysis, we calculate the sum of the attention scores assigned to 2k important tokens selected in each layer and compare this sum to the attention scores allocated to all 128k tokens.
-Despite important tokens comprising only 1.56% of the total number of tokens, they receive a majority of the attention scores.
-This is particularly notable in the later layers, where the proportion of attention scores attributed to these tokens consistently exceeds 50% and often approaches 80%.
-This result indicates that they are central to the LLM processing and the remaining tokens could be removed with a marginal impact on the LLM output.
-3.4 Dynamics of Token Importance
-Figure 4: Minimum match rate between indices of 2k important tokens selected at each layer and its subsequent layers.
-Figure 5: Visualization of output logits by t-SNE for (left) LLaMA-3.1-8B-Instruct and (right) Mistral-Nemo-12B-Instruct.
-Given that tokens discarded after the TSP layer are not propagated in subsequent layers, there must be consistency in identifying important tokens across layers to facilitate TSP.
-This section examines the dynamics of token importance across layers of LLMs.
-Specifically, we investigate whether tokens deemed important based on their high attention scores remain consistent across LLM layers.
-A significant overlap in the set of important tokens would suggest the potential for pre-selecting these tokens before propagating them to subsequent layers.
-To evaluate the consistency of important tokens across layers, we analyze the important tokens in each layer of the LLM and calculate the minimum match rate of the indices of these tokens with those in subsequent layers.
-The minimum match rate serves as an indicator of an information bottleneck.
-Figure 4 presents our findings from this analysis on various LLMs, where each layer extracts 2k important tokens from 128k tokens.
-The minimum match rate in the first layer is below 10%, indicating low correlation between important tokens in the first and later layers.
-However, as we progress to subsequent layers, this match rate increases, surpassing 25% by the middle layer.
-This indicates that among the 2k tokens, at least 512 consistently appear as important across layers.
-Therefore, if the KV cache budget per layer is set to 512, it suffices to examine the 2k important tokens selected from the middle layer.
-The observed consistency in important token selection in later layers likely stems from the increasing robustness of contextual information extraction.
-Each layer in an LLM uses attention scores to integrate token information, progressively enhancing contextual understanding.
-As the model advances through its layers, contextual analysis becomes more refined, enabling a more precise identification of relevant tokens.
-Thus, later layers of LLMs exhibit a consistent recognition of important tokens, underscoring their refined understanding of token importance in the context of the overall sequence.
-3.5 Overall Impact of TSP on LLM Output
-Figure 6: LongBench results of LLaMA-3.1-8B-insturct with propagation of tokens selected at each layer.
-Building on the insights from Section 3.3 and 3.4, we can expect that the removal of unimportant tokens identified from the middle layer will not significantly impact the performance of the LLM for two main reasons.
-First, only a small set of important tokens plays an important role in the attention mechanism and there is the consistency in the identification of important tokens in the middle layer and subsequent layers.
-We analyze real changes in the output logits of an LLM after selecting 2k tokens from each layer and discarding the unselected tokens before propagating them to subsequent layers.
-For this analysis, we use t-SNE (t-distributed Stochastic Neighbor Embedding) to project the final output logits of LLMs into a 2D space.
-Figure 5 demonstrates the distances between output logits obtained with full-context propagation and those obtained through propagation focusing exclusively on important tokens.
-This visual representation shows that removing tokens in the early layers causes significant deviations in output logits compared to full-context propagation.
-However, as expected, removing tokens from later layers preserves a high degree of similarity to the results from full-context models.
-This contributes to stable accuracy results on the LongBench benchmark, as illustrated in Figure 6.
-In light of this observation, we designate the middle layer of each LLM as the TSP layer for the remainder of this paper.
-Moreover, in Figure 6, we compared the accuracy obtained with TSP and a GemFilter-like propagation.
-TSP consistently outperforms GemFilter in terms of accuracy.
-This superior performance can primarily be attributed to how TSP manages the variability in token importance across layers.
-Unlike GemFilter, which reverts to the first layer of the LLM with selected tokens and recomputes the prefill stage, TSP directly propagates the selected tokens to subsequent layers without reverting to the initial layer.
-This direct propagation allows TSP to better accommodate the variability in token importance.
-Furthermore, the large accuracy gap between TSP and GemFilter when tokens are selected in the initial layers highlights the importance of leveraging token information that has been processed with the full context before making important token selections.
-3.6 TSP Length and KV Budget
-The match rate of the important tokens is around 25% (Figure 4).
-This rate is sufficiently high to justify the use of TSP approach with an adequate TSP length (e.g.
-2k).
-However, a short TSP length (e.g.
-512) may lead to significant information loss.
-Therefore, TSP length, which defines the number of tokens selected for propagation, is strategically designed to operate independently of the KV cache budget.
-For instance, if the KV cache budget is set to 512 while the TSP length is 2048, TSP propagates 2048 tokens during its operation, but only the final 512 tokens are stored in the KV cache.
-This separation allows TSP to process a significantly larger set of tokens during propagation, enabling it to capture more contextual information while maintaining the KV cache within its predefined budget.
-As shown in Figure 7, a TSP length of 2048 has been found sufficient to preserve accuracy of the model under specific KV budget while maintaining efficiency in terms of TTFT.
-Table 1: LongBench results comparison between baseline methods and FastKV.
-Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
-Method
-NrtvQA
-Qasper
-MF-en
-HotpotQA
-2WikiMQA
-Musique
-GovReport
-QMSum
-MultiNews
-TREC
-TriviaQA
-SAMSum
-LCC
-RB-P
-Avg.
-LLaMA-3.1-8B-Inustruct, KV Budget = Full
-Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 48.63
+A.2 Comparison with GemFilter
+We compare FastKV with GemFilter, a recent method that accelerates the prefill stage by pruning the input prompt based on token relevance.
+As shown in Figure 9, GemFilter selects a subset of tokens (e.g., 512) at an intermediate filter layer and then restarts the prefill computation using only these tokens.
+This two-stage process, consisting of preprocessing followed by recomputation, enables latency reduction in the prefill stage and effective KV cache compression under a fixed budget.
+However, GemFilter’s strategy of pruning the input prompt leads to an irreversible loss of information from discarded tokens.
+Since only the selected tokens are used for recomputation, the hidden states of the removed tokens are not propagated to the later layers, fundamentally limiting the model’s ability to retain global context.
+As illustrated in Figure 10, this disrupts the full-sequence attention patterns essential for deep semantic understanding.
+As a result, GemFilter shows poor performance on tasks that require comprehensive context, such as code completion.
+In contrast, FastKV mitigates this issue by introducing Token-Selective Propagation (TSP).
+Rather than pruning the input prompt, FastKV allows the early layers to process the full sequence and selects a subset of semantically important tokens (e.g., 2K) at a mid-layer (TSP layer) for propagation.
+The hidden states of all tokens are preserved and propagated until the TSP layer, ensuring that downstream layers receive semantically rich representations.
+This design enables FastKV to preserve contextual fidelity while still achieving significant latency and memory savings.
+Figure 9: Comparison of processing flows between GemFilter and FastKV.
+GemFilter prunes the input prompt and recomputes later layers with only selected tokens, while FastKV propagates the full-context to a mid-layer before reducing the token set for further computation.
+Figure 10: Visualization of attention computation.
+In GemFilter, discarded tokens are never represented in the attention output, leading to fragmented input understanding.
+FastKV retains all tokens through early layers and compresses only after contextual aggregation, preserving global semantics.
+Appendix B Details on Evaluation
+B.1 Models
+We conduct our main evaluation using two open-source LLMs: LLaMA-3.1-8B-UltraLong-1M-Instruct and Mistral-Nemo-12B-Instruct.
+Additionally, we include LLaMA-3.1-8B-Instruct, the base model from which LLaMA-3.1-8B-UltraLong-1M-Instruct is further trained, for supplementary measurements.
+LLaMA-3.1-8B-Instruct supports a context window of 128K tokens.
+The checkpoints for all models are publicly available at:
+B.2 Datasets
+In this section, we provide detailed information on the benchmark datasets used to evaluate FastKV’s performance against baseline methods.
+LongBench.
+LongBench is a collection of benchmark datasets designed to assess the long-context understanding capabilities of LLMs.
+It includes tasks in both English and Chinese, spanning six categories, each consisting of multiple subtasks.
+In this study, we evaluate FastKV on a subset of LongBench that includes only the English-language tasks and code-related benchmarks.
+Detailed information for each selected subtask is provided in Table 3.
+Table 3: LongBench dataset details.
+Task Type Task Abbreviation Eval metric Language #Sample
+Single-doc QA NarrativeQA NrtvQA F1 EN 200
+Qasper Qasper F1 EN 200
+MultiFieldQA-en MF-en F1 EN 150
+HotpotQA HotpotQA F1 EN 200
+Multi-doc QA 2WikiMultihopQA 2WikiMultihopQA F1 EN 200
+MuSiQue MuSiQue F1 EN 200
+Summarization GovReport GovReport Rouge-L EN 200
+QMSum QMSum Rouge-L EN 200
+MultiNews MultiNews Rouge-L EN 200
+Few shot TREC TREC Accuracy EN 200
+TriviaQA TriviaQA F1 EN 200
+SAMSum SAMSum Rouge-L EN 200
+Code LCC LCC Edit Sim Python/C#/Java 500
+RepoBench-P RB-P Edit Sim Python/Java 500
+Synthetic PassageCount PCount Accuracy EN 200
+PassageRetrieval-en PRe Accuracy EN 200
+Needle-in-a-Haystack.
+Needle-in-a-Haystack is a benchmark designed to evaluate the in-context retrieval capabilities of LLMs under long-context settings.
+A statement, referred to as the needle, is randomly selected and inserted at a specific position within a long input context.
+The model is tasked with retrieving the needle content given the entire context.
+Evaluation is conducted across various context lengths and needle depths (insertion positions) to systematically measure retrieval performance under different levels of difficulty.
+In our experiments, retrieval performance is evaluated at context lengths ranging from 32K to 256K tokens, with measurements taken at 32K-token intervals.
+Appendix C More Experimental Results
+C.1 LongBench
+We provide additional LongBench results for LLaMA-3.1-8B-UltraLong-1M-Instruct and Mistral-Nemo-12B-instruct with a KV budget of 1024 (Table 4), as well as a full breakdown of the LongBench results of an additional model, LLaMA-3.1-8B-Instruct (Table 5).
+LLaMA-3.1-8B-Instruct shares the same architecture as LLaMA-3.1-8B-UltraLong-1M-Instruct and serves as its base model, thus exhibiting similar performance trends to the two LLMs reported in Section 4.2.
+Table 4: LongBench results comparison between baseline methods and FastKV with KV budget 1024.
+Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Synthetic
+Method NrtvQA Qasper MF-en HotpotQA 2WikiMQA MuSiQue GovReport QMSum MultiNews TREC TriviaQA SAMSum LCC RB-P PCount PRe Avg.
+LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = Full
+Full KV 23.59 39.33 51.80 56.89 45.53 34.66 33.57 22.96 26.10 73.50 90.88 42.49 62.03 67.05 2.00 89.50 47.62
+LLaMA-3.1-8B-UltraLong-1M-Instruct, KV Budget = 1024
+SnapKV 23.82 37.46 51.63 56.66 45.65 34.55 29.13 22.78 25.31 73.50 91.09 41.65 62.08 65.94 2.00 89.00 47.02
+AdaKV 23.46 37.49 51.54 56.20 44.98 34.91 29.19 22.76 25.31 73.50 91.06 42.37 62.18 67.40 2.00 88.50 47.05
+HeadKV 23.60 37.97 52.03 55.99 45.14 34.92 31.15 23.44 26.12 73.50 90.88 41.28 62.61 67.15 2.00 89.00 47.30
+GemFilter 17.68 29.73 52.20 53.25 49.03 28.34 29.73 20.46 24.46 67.50 88.84 38.43 43.08 45.89 1.00 74.00 41.48
+FastKV 24.12 37.51 52.93 56.65 46.83 34.12 29.09 22.60 25.19 75.00 91.06 42.17 61.96 66.91 2.00 87.50 47.23
+Mistral-Nemo-12B-Instruct, KV Budget = Full
+Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 1.50 98.00 48.51
+Mistral-Nemo-12B-Instruct, KV Budget = 1024
+SnapKV 23.27 41.42 56.01 47.85 46.09 25.83 25.97 23.70 25.37 74.00 89.82 43.48 68.46 65.84 0.50 97.50 47.19
+AdaKV 24.29 41.50 55.45 47.70 46.72 26.12 24.94 22.96 24.84 74.00 89.82 44.11 68.05 65.29 1.00 98.00 47.77
+HeadKV 25.94 41.38 56.45 48.22 45.90 25.05 28.67 23.72 25.72 75.00 89.84 42.64 68.16 66.60 1.50 98.50 47.71
+GemFilter 26.94 40.67 53.63 55.82 56.90 38.23 28.68 20.76 25.34 69.00 87.32 42.49 44.33 45.41 4.50 87.00 45.44
+FastKV 26.26 42.37 57.71 49.40 46.88 25.58 26.17 23.08 25.29 75.00 89.72 43.56 70.13 67.39 0.50 99.00 48.00
+Table 5: LLaMA-3.1-8B-Instruct LongBench results comparison between baseline methods and FastKV.
+Single-Document QA Multi-Document QA Summarization Few-shot Learning Code Synthetic
+Method NrtvQA Qasper MF-en HotpotQA 2WikiMQA MuSiQue GovReport QMSum MultiNews TREC TriviaQA SAMSum LCC RB-P PCount PRe Avg.
+LLaMA-3.1-8B-Instruct, KV Budget = Full
+Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 8.91 99.50 49.33
 LLaMA-3.1-8B-Instruct, KV Budget = 512
-SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 46.60
-AdaKV 29.07 40.16 52.44 53.90 43.05 31.10 25.75 24.39 24.85 69.00 92.34 42.05 63.43 55.32 46.20
-HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 47.76
-GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 35.64
-FastKV 30.38 41.12 55.71 54.35 46.69 30.89 26.78 23.82 24.37 72.50 92.04 42.64 62.60 52.74 46.90
+SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 8.62 99.50 47.54
+AdaKV 30.16 40.11 54.13 54.27 44.73 31.19 27.08 23.82 24.39 71.00 91.90 41.64 60.54 53.15 8.12 99.50 47.23
+HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 8.33 99.50 48.53
+GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 3.53 52.00 34.65
+FastKV 30.28 41.08 55.50 54.35 46.69 30.89 26.81 24.01 24.38 72.50 92.04 42.67 63.99 54.94 7.06 99.50 47.92
+LLaMA-3.1-8B-Instruct, KV Budget = 1024
+SnapKV 31.23 42.52 53.96 55.48 45.43 31.50 29.54 24.78 26.17 70.50 91.73 42.52 62.25 54.72 7.96 99.50 48.11
+AdaKV 30.83 44.33 54.54 55.86 45.39 31.96 28.70 24.67 26.10 71.50 91.48 42.40 61.71 54.81 8.06 99.50 48.24
+HeadKV 30.55 44.66 54.69 55.47 46.20 31.63 32.88 25.14 27.10 73.00 91.57 42.88 63.76 56.06 8.36 99.50 48.97
+GemFilter 19.34 34.76 46.58 46.82 42.83 27.51 29.89 18.96 25.68 63.00 90.70 42.50 38.09 35.14 3.82 86.50 40.76
+FastKV 30.25 44.01 54.72 54.78 46.56 30.59 29.38 24.54 26.23 73.50 91.76 42.58 63.99 57.32 6.86 99.50 48.54
 LLaMA-3.1-8B-Instruct, KV Budget = 2048
-SnapKV 31.31 44.96 55.15 55.48 45.29 30.78 31.92 24.63 26.97 71.50 91.48 43.38 63.14 56.04 48.00
-AdaKV 30.64 45.23 55.46 55.55 45.26 31.14 31.28 24.89 26.72 72.50 91.64 42.75 64.88 59.25 48.37
-HeadKV 30.53 44.99 54.98 55.47 46.08 31.42 32.41 25.05 27.20 72.50 91.57 42.84 63.45 56.75 48.23
-GemFilter 23.64 41.17 51.39 53.97 45.32 29.24 32.01 20.50 26.91 70.00 91.59 42.59 47.35 38.91 43.90
-FastKV 30.31 45.41 54.79 55.11 46.57 30.46 31.25 24.82 27.02 73.50 91.48 43.66 63.17 55.82 48.10
-Mistral-Nemo-12B-Instruct, KV Budget = Full
-Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 48.33
-Mistral-Nemo-12B-Instruct, KV Budget = 512
-SnapKV 23.58 40.12 55.17 47.91 45.69 25.17 23.77 22.48 24.08 74.00 89.44 43.09 68.31 62.15 46.07
-AdaKV 24.46 41.52 57.34 47.89 45.83 25.57 23.58 22.11 23.88 74.00 89.52 43.20 68.21 64.15 46.52
-HeadKV 25.88 40.57 57.12 48.11 46.11 25.26 26.19 22.92 25.25 74.50 89.74 41.81 68.79 64.76 46.93
-GemFilter 25.45 37.64 53.91 52.83 51.26 33.02 26.23 19.42 23.87 65.50 84.16 40.05 38.03 41.13 42.32
-FastKV 25.51 41.48 56.76 48.73 45.43 26.53 23.42 22.15 24.02 74.00 89.42 43.22 68.61 63.57 46.63
-Mistral-Nemo-12B-Instruct, KV Budget = 2048
-SnapKV 24.35 42.95 56.92 49.08 46.43 26.26 28.27 23.68 26.07 75.00 89.82 44.20 68.89 68.00 47.85
-AdaKV 26.02 43.08 57.43 48.76 46.53 26.12 27.80 23.44 25.93 75.00 89.66 44.49 68.53 67.57 47.88
-HeadKV 26.05 42.15 57.46 48.49 45.85 25.86 30.27 24.22 26.17 75.00 89.66 43.74 68.64 67.29 47.92
-GemFilter 26.42 42.40 56.98 57.64 53.92 34.33 30.51 21.60 25.96 72.00 89.65 44.48 48.34 48.06 46.59
-FastKV 26.61 43.37 57.12 49.41 47.22 26.27 28.40 23.22 25.81 75.00 89.72 43.04 68.67 67.61 48.11
-3.7 GQA Compatible KV Cache Compression
-While FastKV incorporates TSP for the latter half of the LLM, it takes a different approach to KV cache compression by propagating full-context data in the early layers and evaluating the importance of KV data using attention scores generated by each layer.
-As the state-of-the-art LLMs Yang et al.
-(2024); Abdin et al.
-(2024) increasingly adopt GQA Ainslie et al.
-(2023), which groups multiple attention heads together and shares KV data within each group, FastKV incorporates GQA compatible KV cache compression technique presented in AdaKV to the early layers.
-This technique adapts SnapKV approach (Equation 1) to GQA by aggregating attention scores across all heads within the group for each token as follows:
-$S_{i}^{l,g}=\frac{1}{H_{G}}\sum_{h=h_{g}}^{h_{g}+H_{G}}S_{i}^{j,h}$ (3)
-Here, $H_{G}$ denotes the number of attention heads in each group.
-$g$ and $h_{g}$ denotes the group index and the index of the first head in the group.
-This aggregation evaluates the importance of tokens at the group level.
-FastKV then selects and stores KV cache entries corresponding to the top-KV budget important tokens in an attention-group-wise manner.
-Figure 7: LongBench results (line) and TTFT (bar) of LLaMA-3.1-8B-insturct across various KV budgets and TSP lengths.
-Figure 8: Needle-in-a-Haystack results of LLaMA-3.1-8B-Instruct with 512 KV budget.
-4 Experiments
-4.1 Setup
-Models & Datasets.
-We evaluate two open-source LLMs of different sizes: LLaMA-3.1-8B-Instruct Dubey et al.
-(2024) and Mistral-Nemo-12B-Instruct Mistral AI Team (2024).
-These models have 32 and 40 decoder layers, respectively, and employ GQA Ainslie et al.
-(2023) with a context window size of 128k tokens.
-To assess long-context understanding and retrieval capabilities, we use two benchmark datasets: LongBench Bai et al. (2023) and Needle-in-a-Haystack Kamradt (2023).
-Implementation Details.
-We integrated our proposed FastKV method upon self-attention implementation of HuggingFace Transformers library, which utilizes FlashAttention-2 Dao (2023) kernel.
-We select layer 15 and 19 as the TSP layer for LLaMA-3.1-8B-Instruct, and Mistral-Nemo-12B-Instruct, respectively.
-The TSP length is set to 2048.
-We fix the observation window size ($N_{obs}$) to 8 and the kernel size of average pooling ($12w_{p}+$ 1) to 7.
-Baselines.
-We compare our approach with four baseline methods for KV cache compression: SnapKV Li et al.
-(2024), AdaKV Feng et al.
-(2024), HeadKV Fu et al.
-(2024), and GemFilter Shi et al.
-(2024).
-SnapKV, AdaKV, and HeadKV use the same local window size, pooling method, and pooling kernel size as FastKV.
-The indices of GemFilter filter layers for LLaMA-3.1-8B-Instruct and Mistral-Nemo-12B-Instruct are 13 and 19, respectively Shi et al.
-(2024).
-4.2 Accuracy Evaluation
-LongBench.
-The accuracy evaluation results on LongBench are summarized in Table 1.
-The full breakdown of results is provided in Appendix A.1.
-Previous works, such as SnapKV, AdaKV, and HeadKV, which use attention-head-wise or attention-group-wise token removal, successfully maintain accuracy after KV cache compression, with an average accuracy drop of less than 2.5% and 0.51% compared to FullKV for 512 and 2048 KV budgets, respectively.
-GemFilter, which discards unselected tokens from the input prompt and recomputes the prefill stage to achieve both memory and computational efficiency, results in a significantly larger accuracy drop, with reductions of up to 13.0% and 4.7% compared to FullKV for 512 and 2048 KV budgets, respectively.
-On the other hand, while the proposed FastKV is designed to achieve memory and computational efficiency comparable to GemFilter, it successfully preserves accuracy on par with SnapKV, AdaKV, and HeadKV.
-It presents the effectiveness of FastKV’s distinct KV cache compression strategies for the early and later layers of LLMs.
-Needle-in-a-Haystack.
-Figure 8 presents the Needle-in-a-Haystack evaluations for LLaMA-3.1-8B-Instruct with 512 KV budget.
-Additional analysis results with various LLM models and KV budgets can be found in Appendix A.2.
-Consistent with the results observed in LongBench, GemFilter shows the worst performance, while FastKV achieves the best performance with a slight improvement over HeadKV.
-Figure 9: (a) TTFT and (b) throughput results across different methods on LLaMA-3.1-8B-Instruct. (dashed line: Full KV)
-4.3 Latency and Throughput Evaluation
-We evaluate the latency of the prefill stage (time-to-first-token (TTFT)) and throughput of token generation with LLaMA-3.1-8B-Instruct model and a single NVIDIA A100 GPU.
-Figure 9 reports TTFT results with 64k/128k input with 512 KV budget, and throughput results with 128k input with 512/2048 KV budget.
-More detailed evaluation results can be found in Appendix A.3.
-TTFT.
-As shown in Figure 9(a), both GemFilter and FastKV achieve significant TTFT improvements over other baselines by reducing the computational complexity of the prefill stage through the concept of propagating only selected tokens during this stage.
-As a result, FastKV achieves TTFT that is 1.60$\times$ shorter and 1.97$\times$ shorter than Full KV, with 64k and 128k input, respecitvely.
-While both methods enhance efficiency, GemFilter incurs additional overhead due to the need for prefill recomputation, whereas FastKV introduces extra processes for compressing generated KV caches.
-These two factors offset each other.
-Consequently, GemFilter’s slight advantage in TTFT is primarily attributed to its earlier filter layer (13) compared to the TSP layer (15) for LLaMA-3.1-8B-Instruct.
-In contrast, SnapKV, AdaKV, and HeadKV exhibit even longer TTFT than Full KV.
-This is because they process the full-context information during the prefill stage, similar to Full KV, while also incurring additional overhead from indexing and selecting crucial tokens.
-Throughput.
-We measure the throughput for generating a total of 128 tokens.
-As shown in Figure 9(b), all KV compression techniques result in throughput improvements compared to Full KV.
-Both FastKV and GemFilter achieve a significant level of throughput improvement.
-FastKV’s throughput reaches 5.07$\times$ of Full KV and 1.40$\times$ of HeadKV with 512 KV budget.
-In contrast, SnapKV, which stores KV caches of different tokens set for each attention head, shows slightly lower throughput.
-Additionally, HeadKV, which allocates varying KV budgets across attention heads, exhibits the lowest throughput among KV cache compression techniques.
-Managing different sequence lengths for keys and values across attention heads increases the complexity of the attention mechanism.
-This effect gets stronger when the KV cache size is large, thereby HeadKV experiences steep decline in throughput as KV budget increases, whereas other methods’ throughput remain relatively consistent.
-AdaKV also adopts fine-grained KV budget allocation similar to HeadKV except that its granularity is attention-group.
-AdaKV mitigates the influence of fragmented KV budget to some extent thanks to GQA-aware compression, but it does not achieve the same level of throughput with FastKV or GemFilter.
-5 Conclusion
-In this paper, we introduce FastKV, a novel KV cache compression method designed to improve the efficiency of long-context LLM inference while preserving accuracy.
-Unlike previous KV cache compression techniques, FastKV enhances computational efficiency by introducing Token-Selective Propagation (TSP).
-This method strategically retains full-context information in early layers and propagates a compressed subset of tokens in later layers, reducing the computational cost of the prefill stage.
-Furthermore, our approach is optimized by integrating GQA compatible KV cache compression, which enables more efficient memory and computation management by leveraging group-wise token selection.
-Our experimental results demonstrate that FastKV achieves higher throughput and lower TTFT than baseline methods while maintaining high accuracy on long-context benchmarks.
-Appendix A More Experimental Results
-A.1 LongBench
-We provide a full breakdown of the LongBench evaluation results for LLaMA-3.1-8B-Instruct (Table 2), LLaMA-3.2-3B-Instruct (Table 3), and Mistral-Nemo-12B-Instruct (Table 4).
-For LLaMA-3.2-3B-Instruct, which has 28 decoder layers, we set the TSP layer and GemFilter filter layer to 13 both.
-Table 2: LongBench results of LLaMA-3.1-8B-Instruct.
-Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
-Method
-NrtvQA
-Qasper
-MF-en
-HotpotQA
-2WikiMQA
-Musique
-GovReport
-QMSum
-MultiNews
-TREC
-TriviaQA
-SAMSum
-LCC
-RB-P
-Avg.
-LLaMA-3.1-8B-Instruct, KV Cache Size = Full
-Full KV 30.21 45.53 55.01 56.01 46.65 31.28 35.13 25.28 27.25 73.00 91.64 43.80 63.38 56.64 48.63
-LLaMA-3.1-8B-Inustruct, KV Budget = 128
-SnapKV 27.39 30.40 50.35 53.04 43.49 28.89 22.87 23.00 21.74 62.00 91.19 40.31 58.43 49.33 43.03
-AdaKV 24.90 24.41 49.95 53.15 41.73 28.55 20.54 23.21 20.28 50.50 89.49 40.71 58.74 52.40 41.33
-HeadKV 29.74 40.34 53.51 54.83 45.35 30.62 26.40 24.23 24.58 71.50 88.51 40.14 60.19 52.92 45.92
-GemFilter 12.50 11.00 16.52 28.84 18.82 13.53 19.66 17.67 14.34 58.17 78.51 32.99 19.92 21.35 25.99
-FastKV 27.63 27.84 52.32 52.00 45.27 28.35 22.00 22.81 20.64 69.50 91.28 40.17 59.08 48.38 42.66
-LLaMA-3.1-8B-Instruct, KV Budget = 512
-SnapKV 30.09 41.62 53.89 54.77 45.12 31.22 27.37 24.16 24.86 71.00 91.90 42.43 61.53 52.49 46.60
-AdaKV 29.07 40.16 52.44 53.90 43.05 31.10 25.75 24.39 24.85 69.00 92.34 42.05 63.43 55.32 46.20
-HeadKV 30.17 44.03 54.29 54.71 46.10 31.58 31.14 24.46 26.66 73.00 91.72 42.17 62.89 55.71 47.76
-GemFilter 18.52 22.15 37.73 46.88 32.56 24.52 27.05 17.75 23.25 60.50 89.49 40.16 27.35 31.02 35.64
-FastKV 30.38 41.12 55.71 54.35 46.69 30.89 26.78 23.82 24.37 72.50 92.04 42.64 62.60 52.74 46.90
-LLaMA-3.1-8B-Instruct, KV Budget = 1024
-SnapKV 31.23 42.52 53.96 55.48 45.43 31.50 29.54 24.78 26.17 70.50 91.73 42.52 62.25 54.72 47.31
-AdaKV 29.23 44.09 53.82 54.80 44.01 31.40 28.86 24.73 26.04 72.50 91.72 42.56 63.22 56.33 47.38
-HeadKV 30.55 44.66 54.69 55.47 46.20 31.63 32.88 25.14 27.10 73.00 91.57 42.88 63.76 56.06 48.26
-GemFilter 19.34 34.76 46.58 46.82 42.83 27.51 29.89 18.96 25.68 63.00 90.70 42.50 38.09 35.14 40.13
-FastKV 30.22 44.01 55.04 54.76 46.76 30.59 29.43 24.66 26.19 73.50 91.76 42.42 62.54 55.02 47.64
-LLaMA-3.1-8B-Instruct, KV Budget = 2048
-SnapKV 31.31 44.96 55.15 55.48 45.29 30.78 31.92 24.63 26.97 71.50 91.48 43.38 63.14 56.04 48.00
-AdaKV 30.64 45.23 55.46 55.55 45.26 31.14 31.28 24.89 26.72 72.50 91.64 42.75 64.88 59.25 48.37
-HeadKV 30.53 44.99 54.98 55.47 46.08 31.42 32.41 25.05 27.20 72.50 91.57 42.84 63.45 56.75 48.23
-GemFilter 23.64 41.17 51.39 53.97 45.32 29.24 32.01 20.50 26.91 70.00 91.59 42.59 47.35 38.91 43.90
-FastKV 30.31 45.41 54.79 55.11 46.57 30.46 31.25 24.82 27.02 73.50 91.48 43.66 63.17 55.82 48.10
-Table 3: LongBench results of LLaMA-3.2-3B-Instruct.
-Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
-Method
-NrtvQA
-Qasper
-MF-en
-HotpotQA
-2WikiMQA
-Musique
-GovReport
-QMSum
-MultiNews
-TREC
-TriviaQA
-SAMSum
-LCC
-RB-P
-Avg.
-LLaMA-3.2-3B-Instruct, KV Budget = Full
-Full KV 23.41 40.60 49.79 48.71 38.89 20.54 34.07 23.65 26.67 71.50 88.89 43.20 52.13 54.16 44.02
-LLaMA-3.2-3B-Instruct, KV Budget= 128
-SnapKV 17.92 23.18 44.59 45.35 35.08 17.07 21.96 21.05 20.38 66.00 87.55 38.39 48.31 45.82 38.05
-AdaKV 19.94 23.15 44.50 45.17 36.51 18.68 21.86 21.46 20.57 67.00 87.82 39.26 49.39 47.61 38.78
-HeadKV 21.59 32.15 48.00 48.79 38.28 18.51 25.30 21.94 23.86 69.50 88.45 38.34 52.11 50.25 41.22
-GemFilter 10.36 10.70 26.36 35.84 27.75 11.63 19.23 17.00 13.59 55.50 71.58 28.40 19.13 27.46 26.75
-FastKV 19.95 23.09 44.34 45.25 34.49 19.89 21.34 20.63 20.02 66.00 88.05 38.37 48.91 47.13 38.39
-LLaMA-3.2-3B-Instruct, KV Budget = 512
-SnapKV 21.03 32.79 49.01 48.68 37.59 20.07 25.91 22.76 24.22 69.00 88.75 41.78 51.68 51.25 41.75
-AdaKV 21.30 32.88 49.92 47.65 37.20 19.72 25.96 22.65 24.31 69.50 89.03 41.55 51.63 51.87 41.80
-HeadKV 22.79 36.45 50.57 48.85 39.16 20.44 29.79 23.24 26.10 70.50 89.66 40.80 53.72 53.83 43.28
-GemFilter 20.88 31.37 44.31 47.60 34.87 23.59 25.89 19.82 24.01 55.00 85.73 35.29 26.68 32.72 36.27
-FastKV 21.73 31.55 49.88 47.31 36.54 22.14 25.47 22.38 24.05 71.50 88.99 40.91 52.46 52.17 41.93
-LLaMA-3.2-3B-Instruct, KV Budget = 1024
-SnapKV 21.71 36.22 49.04 48.78 38.14 21.21 28.32 23.25 25.62 69.00 89.03 41.68 53.01 53.42 42.75
-AdaKV 21.69 36.91 49.76 48.10 39.27 21.25 28.01 22.78 25.66 70.00 88.96 42.62 53.00 53.35 42.96
-HeadKV 23.44 38.06 50.37 49.14 39.51 20.33 31.48 23.26 26.59 71.00 89.53 41.41 53.84 54.14 43.72
-GemFilter 21.34 33.72 48.52 50.13 33.75 20.82 28.90 20.73 24.73 61.50 89.17 38.44 29.89 34.05 38.26
-FastKV 23.13 35.27 49.48 47.93 38.74 21.42 28.09 22.66 25.86 71.50 89.46 41.46 53.09 54.37 43.03
-LLaMA-3.2-3B-Instruct, KV Budget = 2048
-SnapKV 22.89 38.70 50.91 48.38 39.29 20.55 30.79 23.17 26.46 70.50 88.89 42.25 52.16 54.03 43.50
-AdaKV 23.43 39.28 50.26 48.61 39.60 20.82 30.39 22.94 26.65 71.00 88.89 42.57 52.33 54.60 43.67
-HeadKV 22.56 39.84 50.40 49.15 38.95 20.25 32.98 23.73 26.71 71.00 89.28 42.74 53.56 55.30 44.03
-GemFilter 17.48 38.31 49.93 50.57 40.29 19.02 30.95 21.50 26.28 64.50 90.07 40.59 37.97 38.88 40.45
-FastKV 24.02 37.99 50.06 47.58 38.76 21.93 30.45 23.24 26.29 72.50 89.22 41.73 52.71 55.52 43.71
-Table 4: LongBench results of Mistral-Nemo-12B-Instruct.
-Single-Doc QA Multi-Doc QA Summarization Few-shot Learning Coding
-Method
-NrtvQA
-Qasper
-MF-en
-HotpotQA
-2WikiMQA
-Musique
-GovReport
-QMSum
-MultiNews
-TREC
-TriviaQA
-SAMSum
-LCC
-RB-P
-Avg.
-Mistral-Nemo-12B-Instruct, Budget = Full
-Full KV 26.27 43.64 58.11 49.34 45.85 26.26 31.31 24.15 26.08 75.00 89.66 44.32 68.58 68.11 48.33
-Mistral-Nemo-12B-Instruct, KV Budget = 128
-SnapKV 21.29 34.24 51.53 48.07 44.55 23.99 19.93 20.73 20.69 69.50 89.52 39.92 62.85 53.08 42.85
-AdaKV 22.29 36.52 52.50 47.28 44.73 21.84 19.15 20.61 20.74 65.50 88.64 39.25 63.79 54.07 42.64
-HeadKV 22.62 38.76 52.54 46.49 45.97 21.80 20.74 21.37 22.10 73.00 89.42 39.16 64.92 56.15 43.93
-GemFilter 15.74 18.30 42.98 41.86 34.20 18.44 21.72 16.56 18.81 57.00 63.02 31.78 25.74 30.62 31.20
-FastKV 22.26 35.09 53.33 49.15 45.63 23.67 19.15 20.80 20.47 65.00 90.06 39.88 63.65 55.68 43.13
-Mistral-Nemo-12B-Instruct, KV Budget = 512
-SnapKV 23.58 40.12 55.17 47.91 45.69 25.17 23.77 22.48 24.08 74.00 89.44 43.09 68.31 62.15 46.07
-AdaKV 24.46 41.52 57.34 47.89 45.83 25.57 23.58 22.11 23.88 74.00 89.52 43.20 68.21 64.15 46.52
-HeadKV 25.88 40.57 57.12 48.11 46.11 25.26 26.19 22.92 25.25 74.50 89.74 41.81 68.79 64.76 46.93
-GemFilter 25.45 37.64 53.91 52.83 51.26 33.02 26.23 19.42 23.87 65.50 84.16 40.05 38.03 41.13 42.32
-FastKV 25.51 41.48 56.76 48.73 45.43 26.53 23.42 22.15 24.02 74.00 89.42 43.22 68.61 63.57 46.63
-Mistral-Nemo-12B-Instruct, KV Budget = 1024
-SnapKV 23.27 41.42 56.01 47.85 46.09 25.83 25.97 23.70 25.37 74.00 89.82 43.48 68.46 65.84 46.94
-AdaKV 25.33 41.82 57.52 48.02 46.48 26.34 25.68 23.12 25.26 74.50 89.82 44.65 68.49 66.21 47.37
-HeadKV 25.94 41.38 56.45 48.22 45.90 25.05 28.67 23.72 25.72 75.00 89.84 42.64 68.66 66.60 47.41
-GemFilter 26.94 40.67 53.63 55.82 56.90 38.23 28.68 20.76 25.34 69.00 87.32 42.49 44.33 45.41 45.39
-FastKV 26.26 42.37 57.71 49.40 46.88 25.58 26.17 23.08 25.29 75.00 89.72 43.56 68.53 65.79 47.52
-Mistral-Nemo-12B-Instruct, KV Budget = 2048
-SnapKV 24.35 42.95 56.92 49.08 46.43 26.26 28.27 23.68 26.07 75.00 89.82 44.20 68.89 68.00 47.85
-AdaKV 26.02 43.08 57.43 48.76 46.53 26.12 27.80 23.44 25.93 75.00 89.66 44.49 68.53 67.57 47.88
-HeadKV 26.05 42.15 57.46 48.49 45.85 25.86 30.27 24.22 26.17 75.00 89.66 43.74 68.64 67.29 47.92
-GemFilter 26.42 42.40 56.98 57.64 53.92 34.33 30.51 21.60 25.96 72.00 89.65 44.48 48.34 48.06 46.59
-FastKV 26.61 43.37 57.12 49.41 47.22 26.27 28.40 23.22 25.81 75.00 89.72 43.04 68.67 67.61 48.11
-A.2 Needle-in-a-Haystack
-We performed further Needle-in-a-Haystack evaluations with varying KV budgets for LLaMA-3.1-8B-Instruct (Figure 10), LLaMA-3.2-3B-Instruct (Figure 11), and Mistral-Nemo-12B-Instruct (Figure 12).
-For LLaMA-3.1-8B-Instruct, the overall retrieval trend remains consistent.
-FastKV continues to achieve strong performance across multiple budgets, often matching or surpassing baseline methods, including HeadKV.
-By contrast, LLaMA-3.2-3B-Instruct produces notably lower scores overall, even when using Full KV.
-Its smaller parameter count appears to limit retrieval capabilities at long contexts, making it more difficult for fine-grained KV allocation techniques to improve performance.
-Although HeadKV excelled with LLaMA-3.1-8B-Instruct, it does not show similar gains on LLaMA-3.2-3B-Instruct, suggesting that certain highly flexible optimizations may not translate effectively to smaller models.
-Nonetheless, FastKV outperforms SnapKV, HeadKV, and GemFilter, and achieves an average score on par with AdaKV, indicating that its core compression strategy retains efficacy when applied to a compact model.
+SnapKV 31.31 44.96 55.15 55.48 45.29 30.78 31.92 24.63 26.97 71.50 91.48 43.38 63.14 56.04 8.21 99.50 48.73
+AdaKV 31.03 45.39 55.22 55.10 45.88 30.31 30.96 24.46 26.93 72.00 91.65 43.76 63.02 56.10 8.18 99.50 48.72
+HeadKV 30.53 44.99 54.98 55.47 46.08 31.42 32.41 25.05 27.20 72.50 91.57 42.84 63.45 56.75 7.99 99.50 48.92
+GemFilter 23.64 41.17 51.39 53.97 45.32 29.24 32.01 20.50 26.91 70.00 91.59 42.59 47.35 38.91 6.88 95.50 44.81
+FastKV 30.38 45.35 54.82 55.11 46.57 30.46 31.16 24.73 26.84 73.50 91.48 43.62 64.84 58.20 6.65 99.50 48.95
+C.2 Needle-in-a-Haystack
+We performed further Needle-in-a-Haystack evaluations with varying KV budgets for LLaMA-3.1-8B-UltraLong-1M-Instruct (Figure 11), and Mistral-Nemo-12B-Instruct (Figure 13).
+Given that Mistral-Nemo-12B-Instruct supports a context window of up to 128K tokens, its evaluation was limited to a context length of 128K.
+For LLaMA-3.1-8B-UltraLong-1M-Instruct, the overall retrieval trend remains consistent.
+FastKV continues to achieve strong performance across multiple budgets, often matching or surpassing baseline methods, including fine-grained KV budget methods such as HeadKV and AdaKV.
 Mistral-Nemo-12B-Instruct exhibits a shorter effective context window than the nominal 128K, frequently struggling to retrieve needles located at extreme depths while handling nearer contexts reasonably well.
-Surprisingly, GemFilter significantly outperform Full KV, and FastKV also records above-Full-KV scores.
-GemFilter discards a subset of tokens and restarts its prefill stage with the reduced sequence, and FastKV continues prefilling the remaining layers with reduced hidden states after TSP.
-This may lead temporary enhancement of retrieval by focusing attention on fewer tokens.
-However, GemFilter’s poor results in LongBench confirm that these gains do not generalize to complex tasks and models.
+For Mistral-Nemo-12B-Instruct, GemFilter significantly outperform FastKV and other baselines.
+GemFilter’s policy that restarts prefilling stage with the reduced token sequence may lead to temporary enhancement of retrieval by focusing attention on fewer tokens.
+However, this trend does not generalize to complex tasks and other models.
+Figure 11: Needle-in-a-Haystack results of LLaMA-3.1-8B-UltraLong-1M-Instruct.
 Figure 12: Needle-in-a-Haystack results of LLaMA-3.1-8B-Instruct.
-Figure 11: Needle-in-a-Haystack results of LLaMA-3.2-3B-Instruct.
 Figure 13: Needle-in-a-Haystack results of Mistral-Nemo-12B-Instruct.
-A.3 Latency and Throughput Evaluation
-We present the TTFT and the throughput results of the baselines and FastKV in Table 5 and Table  7.
-The throughput results of Full KV is depicted in Table  6 for effective comparison.
-The throughput results of KV cache compression methods are evaluated with input context length 128k.
-As the input context length increases, the relative TTFT gain of FastKV compared to SnapKV, AdaKV and HeadKV increases.
-For Mistral-Nemo-12B-instruct, FastKV achieves TTFT similar to GemFilter because the TSP layer index is same with the GemFilter filter layer.
-The token throughput gain of the baselines and FastKV using fixed KV budget grows rapidly as the throughput of Full KV declines due to the increasing size of KV cache.
+C.3 Latency and Throughput Evaluation
+We present the TTFT and throughput results of baseline methods and FastKV in Table 6 and Table 8.
+The TTFT of LLaMA-3.1-8B-UltraLong-1M-Instruct is measured with input context lengths up to 256K tokens, while for Mistral-Nemo-12B-Instruct, the measurement is limited to 128K tokens due to its 128K context window size.
+The throughput results for Full KV are provided separately in Table 7 for effective comparison.
+The throughput results for KV cache compression methods are evaluated at an input context length of 128K tokens.
+As the input context length increases, the relative TTFT gain of FastKV over SnapKV, AdaKV, and HeadKV becomes more pronounced.
+For Mistral-Nemo-12B-Instruct, FastKV achieves TTFT comparable to GemFilter, as both utilize the same TSP layer structure as the GemFilter filtering layer.
+The token throughput of the baselines and FastKV with a fixed KV budget improves rapidly as the throughput of Full KV declines with increasing KV cache size.
 FastKV consistently achieves the highest throughput across all scenarios.
@@
       Context Length
-Method 8k 16k 32k 64k 128k
-LLaMA-3.1-8B-Instruct
-     Full KV 0.64 1.41 3.40 9.05 27.94
-     SnapKV 0.67 1.45 3.49 9.22 28.27
-     AdaKV 0.75 1.53 3.57 9.31 28.36
-     HeadKV 0.70 1.48 3.51 9.25 28.30
-     GemFilter 0.44 0.78 1.66 4.15 12.51
-     FastKV 0.41 0.80 1.81 4.66 14.15
+      Method       8K       16K       32K       64K       128K       256K
+      LLaMA-3.1-8B-UltraLong-1M-Instruct
+      Full KV       0.64       1.41       3.40       9.05       29.2       OOM
+      SnapKV       0.67       1.45       3.49       9.22       29.68       102.60
+      AdaKV       0.75       1.53       3.57       9.31       29.88       103.12
+      HeadKV       0.70       1.48       3.51       9.25       29.65       102.71
+      GemFilter       0.44       0.80       1.66       4.15       12.95       45.08
+      FastKV       0.41       0.80       1.81       4.66       14.80       51.68
       Mistral-Nemo-12B-Instruct
-     Full KV 0.92 2.04 4.85 12.71 37.94
-     SnapKV 0.95 2.08 4.90 12.80 38.11
-     AdaKV 1.05 2.18 5.01 12.92 38.23
-     HeadKV 0.99 2.12 4.95 12.85 38.18
-     GemFilter 0.69 1.25 2.65 6.59 19.22
-     FastKV 0.59 1.16 2.57 6.53 19.22
+      Full KV       0.92       2.04       4.85       12.71       37.94       -
+      SnapKV       0.95       2.08       4.90       12.80       38.11       -
+      AdaKV       1.05       2.18       5.01       12.92       38.23       -
+      HeadKV       1.0       2.1       5.0       12.9       38.18       -
+      GemFilter       0.69       1.25       2.65       6.59       19.22       -
+      FastKV       0.59       1.16       2.57       6.53       19.22       -
 Table 7: Throughput (Tokens/sec) with full KV budget at varying context length.
     Context Length 8K 16K 32K 64K 128K
-    LLaMA-3.1-8B-Instruct 51.55 43.07 32.29 21.24 12.60
+    LLaMA-3.1-8B-UltraLong-1M-Instruct 51.5 43.07 32.29 21.24 12.60
     Mistral-Nemo-12B-Instruct 36.10 30.88 23.67 16.07 9.68
@@
             512             1024             2048
-        LLaMA-3.1-8B-Instruct
+            LLaMA-3.1-8B-UltraLong-1M-Instruct
             SnapKV             48.71             48.62             48.29